### v7

# Cohomology of Lie algebroids over Algebraic spaces

Abstract.
We consider Lie algebroids over an algebraic space (or topological ringed space) as quasicoherent sheaves of Lie-Rinehart algebras. We express hypercohomology for a locally free Lie algebroid (not necessarily of finite rank) as a derived functor, and simplify it via Čech cohomology. Furthermore, we define the Hochschild hypercohomology of a sheaf of generalized bialgebras and study the cases of the universal enveloping algebroid and the jet algebroid of a Lie algebroid. In the sequel, we present a version of Hochschild-Kostant-Rosenberg theorem for a locally free Lie algebroid, as well as its dual version.
Key words and phrases:
Lie algebroids, Lie-Rinehart algebras, tangent sheaf, universal enveloping algebra, Lie algebroid cohomology, bialgebras
AMS Mathematics Subject Classification : $16$S30303030, $16$T10,1010,10 ,$17$B56565656.
1. Introduction
The notion of Lie algebroids plays a prominent role in differential geometry and mathematical physics as they represent generalized infinitesimal symmetries of spaces, which are related to the corresponding global symmetries of spaces described by Lie groupoids [21]. Lie algebroids over a $C^{\infty}$-manifold are joint generalization of tangent vector bundle over the manifold and Lie algebras [26]. An algebraic analogue of Lie algebroids, known as Lie-Rinehart algebras, is used to study general situations [16, 18, 22, 3]. In the context of complex geometry and algebraic geometry, Lie algebroids over analytic spaces [32] and over algebraic varieties (or schemes) [18, 7, 3] respectively have been studied in sheaf theoretic language, which are joint generalization of the tangent sheaf and sheaves of Lie algebras [18, 33, 32]. We study Lie algebroids over an algebraic space (shortly, $a$-space) or over a topological ringed space $(X,\mathcal{O}_{X})$ as certain quasicoherent sheaves of Lie-Rinehart algebras [37, 29], where $\mathcal{O}_{X}$ is a sub-sheaf of algebras of the sheaf of continuous functions $C_{X}^{0}$ on $X$ [33, 30]. This framework unifies the concept of Lie algebroids across these three types of base spaces as special cases. To study calculus on these (smooth and singular) geometric objects in a consistent manner, we need the notion of Lie algebroids within algebro-geometric settings. It allows one to treat several geometric structures, such as Poisson analytic spaces [32], singular foliations or generalized involutive distributions [10, 32, 20], as well as free Lie algebroids [18]. In particular, the sheaf of logarithmic derivations for some (principal or free) divisor of a complex manifold (or a smooth algebraic variety) [4, 32, 29], and the path algebroid of a smooth manifold [18], are key object of study in this context.
Associated with a Lie algebroid over an $a$-space, there are two canonical sheaves of generalized bialgebras: the universal enveloping algebroid and the jet algebroid [18, 7, 6, 2, 32, 29]. Both play crucial roles in studying the homological algebra of a Lie algebroid. Here, the sheaves of generalized bialgebras serve as a sheaf-theoretic analogue (or global version) of the notion of $R/{\mathbb{K}}$-bialgebras [26] or left $R$-bialgebroids [19], referred to as $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras [29]. The universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ of a Lie algebroid $\mathcal{L}$ generalizes the notion of sheaf of differential operators on a manifold [35, 29]. It is sheafification of the presheaf of universal enveloping algebras of Lie-Rinehart algebras associated with each space of sections of $\mathcal{L}$ [3, 29]. It has a canonical $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra structure [29], similar structures are present in [18, 7] by different names. Moreover the dual of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, $\mathscr{J}(\mathcal{O}_{X},\mathcal{L}):=\mathscr{H}om_{\mathcal{O}_{X}}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}),\mathcal{O}_{X})$ is the jet algebroid of $\mathcal{L}$, generalizes the notion of sheaf of jets on a manifold (see[18, 7, 5, 32, 2]). It is sheafification of the presheaf of jet algebras associated with the sheaf of Lie-Rinehart algebras [2, 29]. It also has a canonical $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra structure induced from the structure of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ [29]. Moreover, an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra gives rise to a Lie algebroid over the $a$-space $(X,\mathcal{O}_{X})$, consisting of its sheaf of primitive elements [29].
Cohomology of a Lie algebroid generalizes both Lie algebra cohomology (Chevalley-Eilenberg cohomology) and de Rham cohomology, and is commonly refereed to as Chevalley-Eilenberg-de Rham cohomology [21, 32, 2, 1, 3]. The cohomology of a Lie algebroid $\mathcal{L}$, which is a locally free $\mathcal{O}_{X}$-module of finite rank, over a Noetherian separated scheme [3] or over a complex manifold $(X,\mathcal{O}_{X})$ [32], is described using the derived functor $Ext$. In non-commutative geometry [17, 13, 35], the dual pair consisting of the Hochschild cohomology and homology $(HH^{\bullet}(A),HH_{\bullet}(A))$ of an associative $\mathbb{K}$-algebra $A$ serves as a non-commutative analogue to the classical dual pair $(\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{T}_{X},\wedge^{\bullet}_{\mathcal{O}_{X}}\Omega^{1}_{X})$, which consists of the sheaf of multivector fields and the sheaf of differential forms over a smooth $a$-space $(X,\mathcal{O}_{X})$. This correspondence is given by the standard Hochschild-Kostant-Rosenberg (shortly, HKR) theorem. In [7, 5, 6], the authors establish a version of the HKR theorem, using the dual pair $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$, to explore precalculus up to homotopy for locally free Lie algebroids of finite rank. For that, they introduce an isomorphism of Gerstenhaber algebras, given by a twisted version of the HKR morphism. In this paper, we deduce analogous cohomological results using a different approach to include more general situations, such as relaxing the finite rank condition. To derive such results, we present a global version of cohomological results on Lie-Rinehart algebras [34, 16, 19, 25].
In Section 2, we recall some of the preliminary notions required for the article. In Section 3, we consider Lie algebroid hypercohomology over certain special $a$-spaces. Here, we show that one can define Lie algebroid cohomology, where the underlying $\mathcal{O}_{X}$-module, while locally free, is not necessarily of finite rank, thus encompassing all three geometric setups. The cohomology of a locally free Lie algebroid $\mathcal{L}$ over an $a$-space $(X,\mathcal{O}_{X})$ with coefficient in a $\mathcal{L}$-module $\mathcal{E}$, is given by $\mathbb{H}^{\bullet}(\mathcal{L},\mathcal{E})\cong Ext^{\bullet}_{\mathscr{U}(\mathcal{O}_{X},\mathcal{L})}(\mathcal{O}_{X},\mathcal{E})$. Here, we use the result of Lie-Rinehart cohomology as derived functor for a projective Lie-Rinehart algebra [34, 16] as a local description. We apply this result for the sheaf of logarithmic derivations and for $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras. Using the cohomology groups we consider (first) Chern classes [24, 25] in this general set up. Furthermore, we express the Lie algebroid cohomology in terms of Čech cohomology by considering a good open cover [33, 36, 1]. In Section 4, we consider algebraic (analytic) de Rham cohomologies for some free divisors associated with principal ideal sheaves [29] and compute the corresponding logarithmic de Rham cohomologies [4, 27]. We view these cohomologies as Lie algebroid (hyper)cohomology as described in [32, 1, 3].
In the first part of Section 5, we define Hochschild hypercohomology of an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra. As special cases of it, we study Hochschild hypercohomology of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ associated with a Lie algebroid $\mathcal{L}$. After that, we present a version of Hochschild-Kostant-Rosenbergh (HKR) theorem for locally free Lie algebroids (locally free $\mathcal{O}_{X}$-modules but not necessarily of finite rank) over any of the special $a$-spaces. It provides an isomorphism of graded vector spaces between Hochschild hypercohomology of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and hypercohomology of the sheaf of $\mathcal{L}$-poly vector fields, that is $\mathbb{H}H^{\bullet}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}))\cong\mathbb{H}^{\bullet}(X,\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L})$. This result we get by locally using the algebraic counterpart for projective Lie-Rinehart algebras described in [19]. We derive a result in sheaf cohomology context to simplify the hypercohomology $\mathbb{H}^{\bullet}(X,\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L})$. Moreover, we present the HKR theorem in some of the special cases. Next we discuss about the dual version of HKR theorem in the generalize setup. Thus we show that for a locally free Lie algebroids $\mathcal{L}$ of finite rank over any of the special $a$-space $(X,\mathcal{O}_{X})$, we get a canonical isomorphism of vector spaces between Hochschild hypercohomology of $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ and the Lie algebroid (hyper)cohomology of $\mathcal{L}$, that is $\mathbb{H}H^{\bullet}(\mathscr{J}(\mathcal{O}_{X},\mathcal{L}))\cong\mathbb{H}^{\bullet}(\mathcal{L},\mathcal{O}_{X})$. This result we get by locally using the algebraic counterpart as described in [19]. Both of these hypercohomologies are obtained by the derived functor $Cotor$. At last, we apply the dual HKR theorem for the tangent sheaf $\mathcal{L}:=\mathcal{T}_{X}$ over non-singular $a$-spaces and obtain some interesting results using the (smooth, analytic or algebraic) de Rham cohomology [8, 36] of $X$.
Acknowledgment. The author acknowledges support from the C. S. I. R. fellowship grant during his PhD at Indian Institute of Technology Kanpur. He wholeheartedly expresses his deepest gratitude to Prof. Ashis Mandal for his immense support and guidance. Additionally, he would like to thank Dr. Satyendra Kumar Mishra for his valuable comments and suggestions, which helped to improve the article.
2. Preliminaries
In this section, we recall some standard notions, namely algebraic spaces (or topological ringed spaces), Lie algebroids, universal enveloping algebroids, generalized bialgebras etc in algebro-geometric language. Then, we recall some important relationships among them. In the later sections, we use these notions to consider cohomology theoretic results for such Lie algebroids.
The notation $\mathbb{K}$ is used for $\mathbb{R,C}$ (the real or complex number fields respectively) or a general algebraically closed field of characteristic zero. We denote the constant sheaf by $\mathbb{K}_{X}$ on a topological space $X$ with stalks being isomorphic to $\mathbb{K}$. The sheaf of $\mathbb{K}$-valued continuous functions on $X$ is denoted by $C_{X}^{0}$.
2.1. Lie algebroids over algebraic spaces
Here we consider some special locally ringed spaces [38], and then define Lie algebroids over these spaces [29].
Definition 2.1.
Let $X$ be a topological space and $\mathcal{O}_{X}$ a $\mathbb{K}_{X}$-subalgebra of the sheaf of continuous functions $C_{X}^{0}$ on $X$. The pair $(X,\mathcal{O}_{X})$ is said to be an algebraic space, or simply, an $a$-space.
For an $a$-space $(X,\mathcal{O}_{X})$, the sheaf of derivations $\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X})$ of $\mathcal{O}_{X}$, known as the tangent sheaf of $X$.
Note 2.2.
Consider smooth manifolds, complex manifolds, analytic spaces and algebraic varieties, all with their associated structure sheaf, as special $a$-spaces. The tangent sheaf over such an $a$-space $(X,\mathcal{O}_{X})$ is a quasicoherent module over $\mathcal{O}_{X}$ [18, 32].
Definition 2.3.
A Lie algebroid $\mathcal{L}$ over an $a$-space $(X,\mathcal{O}_{X})$ is a quasicoherent sheaf of $(\mathbb{K}_{X},\mathcal{O}_{X})$-Lie-Rinehart algebras. That is, $\mathcal{L}$ is a $\mathbb{K}_{X}$-Lie algebra and a quasicoherent $\mathcal{O}_{X}$-module equipped with a homomorphism $\mathfrak{a}:(\mathcal{L},[\cdot,\cdot])\rightarrow(\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X}),[\cdot,\cdot]_{c})$ of $\mathcal{O}_{X}$-modules and $\mathbb{K}_{X}$-Lie algebras, called the anchor map. The map $\mathfrak{a}$ satisfies the Leibniz rule: $[D,fD^{\prime}]=f[D,D^{\prime}]+\mathfrak{a}(D)(f)D^{\prime}$ for $f\in\mathcal{O}_{X}$ and $D,D^{\prime}\in\mathcal{L}$.
We denote this Lie algebroid as $(\mathcal{L},[\cdot,\cdot],\mathfrak{a})$ or simply by $\mathcal{L}$.
Remark 2.4.
Lie algebroids over a smooth manifold $X$ is equivalent to locally free Lie algebroids of finite rank over the $a$-space $(X,C_{X}^{\infty})$, where $C_{X}^{\infty}$ is the sheaf of $C^{\infty}$-functions on $X$.
A homomorphism of Lie algebroids over an $a$-space
$\phi:(\mathcal{L}_{1},[\cdot,\cdot]_{1},\mathfrak{a}_{1})\rightarrow(\mathcal{L}_{2},[\cdot,\cdot]_{2},\mathfrak{a}_{2})$
is a sheaf homomorphism of Lie-Rinehart algebras.
Example 2.5.
The standard Lie algebroid structure on the tangent sheaf of an $a$-space $(X,\mathcal{O}_{X})$ is described by the pair $(\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X}),[\cdot,\cdot]_{c},id)=:\mathcal{T}_{X}$. In particular, when $X$ is a real smooth manifold (complex manifold), the sheaf of smooth (holomorphic) vector fields $\mathfrak{X}_{X}$ on $X$ is isomorphic to $\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X})$, forming a locally free Lie algebroid of rank equals to dimension of the manifold $X$.
The sheaf of differential $1$-forms $\Omega^{1}_{X}$ over a Poisson manifold $X$ has a canonical Lie algebroid structure. For smooth and analytic cases the details can be found in [10] and [32] respectively.
Example 2.6.
A singular foliation $\mathcal{F}$ on a real smooth manifold or a complex manifold $(X,\mathcal{O}_{X})$ is an $\mathcal{O}_{X}$-submodule of the Lie algebroid $\mathcal{T}_{X}$ ((((or $\mathfrak{X}_{X})$, which is $(i)$ stable under the Lie bracket and $(ii)$ locally finitely generated. It provides a Stefan-Sussmann distribution on $X$, forms a Lie algebroid over $(X,\mathcal{O}_{X})$. In particular, regular foliations arise from involutive sub bundles of the tangent bundle ((((providing Frobenius distributions)))). See [10, 32, 20] for details.
Example 2.7.
Let $X$ be a complex manifold and $\mathcal{O}_{X}$ be the sheaf of holomorphic functions. The vanishing set ((((or the zero locus)))) $Y:=V(\mathcal{I})$ of a ideal-sheaf $\mathcal{I}\subset\mathcal{O}_{X}$ is a subspace of $X$ which is not necessarily a submanifold ((((it may have singularity as well)))). The sheaf of functions $\mathcal{O}_{Y}:=\mathcal{O}_{X}/{\mathcal{I}}$ on $Y$ is its structure sheaf and the pair $(Y,\mathcal{O}_{Y})$ is an analytic space (see in [32]).
Here, the tangent sheaf $\mathcal{T}_{Y}:=(\mathcal{D}er_{\mathbb{C}_{Y}}(\mathcal{O}_{Y}),[\cdot,\cdot]_{c},id_{Y})$ is not necessarily a locally free Lie algebroid over $(Y,\mathcal{O}_{Y})$. Consider the sheaf of logarithmic derivations ((((forms a generalized involutive distribution or singular foliation)))) [32, 29] as
$\mathcal{T}_{X}(-logY):=\{D\in\mathcal{T}_{X}:D(\mathcal{I})\subset\mathcal{I}\}\hookrightarrow\mathcal{T}_{X}$
((((geometrically, it represents the sheaf of vector fields on $X$ that are tangent to $Y$ for a smooth divisor $Y)$ with the canonical Lie algebroid structure. It is associated with $\mathcal{T}_{Y}$ via the Lie algebroid epimorphism
$\rho:\mathcal{T}_{X}(-logY)\rightarrow\mathcal{T}_{Y}$
defined by $\rho(D)=\bar{D}$, $\bar{D}([{f}])=[{D(f)}]$ for any sections $f\in\mathcal{O}_{X},D\in\mathcal{T}_{X}(-logY)$.
For example, let $X=\mathbb{C}^{2}$ with global coordinate functions $x,y$ and consider the analytic space as $(Y,\mathcal{O}_{Y})$ associated to the vanishing set of the ideal sheaf $\mathcal{I}$ generated by the function $xy$. Here, the subspace $Y$ is the union of points on the axes, known as a normal crossing divisor, has a singularity at the origin $\bar{0}:=(0,0)$. Thus, the Lie algebroid $\mathcal{T}_{X}(-logY)$ is a locally free $\mathcal{O}_{X}$-module generated by the derivations $x\partial_{x}$ and $y\partial_{y}$, but the Lie algebroid $\mathcal{T}_{Y}$ is not locally free. See [32, 29] for details.
2.2. Universal enveloping algebroid of a Lie algebroid
Let $(\mathcal{L},[\cdot,\cdot],\mathfrak{a})$ be a Lie algebroid over any of the special $a$-space $(X,\mathcal{O}_{X})$. For each open set $U$ of $X$, we find the universal enveloping algebra $\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ of the $(\mathbb{K},\mathcal{O}_{X}(U))$-Lie-Rinehart algebra $\mathcal{L}(U)$ or of the Lie-Rinehart pair $(\mathcal{O}_{X}(U),\mathcal{L}(U))$ (see [34, 16, 26]). The sheafification of the cannonical presheaf:
$U\mapsto\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U)),$
is known as the universal enveloping algebroid of the Lie algebroid $\mathcal{L}$, and denoted by $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ (see [5, 32, 3, 35, 29] for details).
From the construction of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, it is an associative $\mathbb{K}_{X}$-algebra and $\mathcal{O}_{X}$-bimodule. Moreover, there is a canonical $\mathbb{K}_{X}$-algebra monomorphism $\iota_{X}:\mathcal{O}_{X}\hookrightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and an $\mathcal{O}_{X}$-linear map $\iota_{\mathcal{L}}:\mathcal{L}\rightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$. The sheaf $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ is generated by $\mathcal{O}_{X}$ and $\iota_{\mathcal{L}}(\mathcal{L})$ satisfy the following identities:
[…]
where $D,D^{\prime}\in\mathcal{L}$,  $f\in\mathcal{O}_{X}$, and $\bar{D}=\iota_{\mathcal{L}}(D)$ for all $D\in\mathcal{L}$.
Hence, the map $\iota_{\mathcal{L}}$ can also be viewed as a $\mathbb{K}_{X}$-Lie algebra homomorphism.
Remark 2.8.
The universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ of $\mathcal{L}$ is characterized by the following universal property:
Let $\mathcal{A}$ be a sheaf of unital associative $\mathbb{K}_{X}$-algebra with sheaf homomorphisms $\phi:\mathcal{O}_{X}\rightarrow\mathcal{A}$ of $\mathbb{K}_{X}$-unital algebras and $\psi:(\mathcal{L},[\cdot,\cdot])\rightarrow(\mathcal{A},[\cdot,\cdot]_{c})$ of $\mathcal{O}_{X}$-linear $\mathbb{K}_{X}$-Lie algebras such that $\phi(f)\psi(D)=\psi(fD)$ and $[\psi(D),\phi(f)]_{c}=\phi(\mathfrak{a}(D(f)))$ holds for $f\in\mathcal{O}_{X}$ and $D\in\mathcal{L}$. Then, there exists a unique $\mathcal{O}_{X}$-linear homomorphism of unital $\mathbb{K}_{X}$-algebras $\widetilde{\psi}:\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathcal{A}$ such that $\widetilde{\psi}\circ\iota_{X}=\phi$ and $\widetilde{\psi}\circ\iota_{\mathcal{L}}=\psi$.
Note 2.9.
For a locally free Lie algebroid $\mathcal{L}$ over an $a$-space $(X,\mathcal{O}_{X})$, the homomorphism $\iota_{\mathcal{L}}:\mathcal{L}\rightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ become an embedding (as $\mathbb{K}_{X}$-algebras and $\mathcal{O}_{X}$-modules).
Note 2.10.
For a non-singular special $a$-space $(X,\mathcal{O}_{X})$, the universal enveloping algebroid is isomorphic to the sheaf of differential operators $\mathcal{D}_{X}$ on $X$ (i.e. the sheaf of differential operators over $\mathcal{O}_{X}$, sometimes denoted as $\mathcal{D}iff(\mathcal{O}_{X}))$, i.e.
$\mathscr{U}(\mathcal{O}_{X},\mathcal{T}_{X})\cong\mathcal{D}iff(\mathcal{O}_{X})=:\mathcal{D}_{X}.$
Note 2.11.
The sheaf of logarithmic derivations and sheaf of logarithmic differential operators are denoted as $\mathcal{T}_{X}(-logY)$ and $\mathcal{D}_{X}(-logY)$ respectively for a principal divisor $Y$ in some complex manifold or smooth algebraic variety $X$ [29].
In the case of a free divisor $Y$ in $X$ (i.e. $\mathcal{T}_{X}(-logY)$ is locally free $\mathcal{O}_{X}$-module [4, 32, 29]) we have (sheafifying the local description given for the module of logarithmic derivations in [22]) the isomorphism
$\mathscr{U}(\mathcal{O}_{X},\mathcal{T}_{X}(-logY))\cong\mathcal{D}_{X}(-logY).$
Remark 2.12.
In [18], the notion of path algebroid $\mathcal{P}_{X}$ of a smooth manifold or smooth algebraic variety $X$ is constructed as the free Lie algebroid generated by tangent sheaf $\mathcal{T}_{X}$. It forms a locally free sheaf of Lie-Rinehart algebras over $(X,\mathcal{O}_{X})$ of infinite rank. The universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{P}_{X})=:\mathbb{D}_{X}$ is described as sheaf of non commutative differential operators on $X$.
2.3. $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras
The notation $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras [29] is considered as a sheaf theoretic analogue of the notion $R/{\mathbb{K}}$-bialgebras [26] or left bialgebroids [19]. As an example, we consider the universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ (and its dual) of a Lie algebroid $\mathcal{L}$ over $(X,\mathcal{O}_{X})$. Similar structure is appeared in [18, 7, 5, 2], in complex and algebraic geometry context.
Definition 2.13.
A tuple $(\mathcal{A},\Delta,\epsilon)$ is called $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra if the following conditions hold.
  • $\mathcal{A}$ is a sheaf of unital associative $\mathbb{K}_{X}$-algebra extending the sheaf $\mathcal{O}_{X}$;
  • $\mathcal{A}$ is equipped with the morphism of sheaf of $\mathcal{O}_{X}$-modules the comultiplication $\Delta:\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathcal{O}_{X}}\mathcal{A}$ and counit $\epsilon:\mathcal{A}\rightarrow\mathcal{O}_{X}$;
  • The image $\Delta(\mathcal{A})$ lies in a certain $\mathbb{K}_{X}$-subalgebra of $\mathcal{A}{\otimes}_{\mathcal{O}_{X}}\mathcal{A}$;
  • $\Delta(1)=1\otimes 1$ and $\epsilon(1)=1$, where $1$ is the unit of $\mathcal{O}_{X}$;
  • $\Delta(ab)=\Delta(a)\Delta(b)$, $\epsilon(ab)=\epsilon(a\epsilon(b))$ for any two sections $a,b$ in $\mathcal{A}$.
We recall the canonical $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra structures on $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and its dual in the following.
The universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ as an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra:
Let $(\mathcal{L},[\cdot,\cdot],\mathfrak{a})$ be a Lie algebroid over one of the special $a$-space $(X,\mathcal{O}_{X})$. The sheaf $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ has a canonical associative unital $\mathbb{K}_{X}$-algebras and an $\mathcal{O}_{X}$-bimodule structure as discussed above.
Observe that $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ is a cocommutative counital $\mathcal{O}_{X}$-coalgebra in the sense that there is a natural cocommutative coassociative comultiplication $\Delta:\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\otimes_{\mathcal{O}_{X}}\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, and counit $\epsilon:\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathcal{O}_{X}$ respectively, which are locally given by the following formulas [5]
$\Delta(f)=f\otimes 1=1\otimes f$,
$\Delta(\bar{D})=\bar{D}\otimes 1+1\otimes\bar{D}$,
$\Delta(D^{\prime}D^{\prime\prime})=\sum D^{\prime}_{(1)}D^{\prime\prime}_{(1)}\otimes D^{\prime}_{(2)}D^{\prime\prime}_{(2)}$,
$\epsilon(\tilde{D})=\tilde{D}(1)$,
for a section $f$ of $\mathcal{O}_{X}$, for a section $D$ of $\mathcal{L}$ with $\bar{D}=\iota_{\mathcal{L}}(D)$, and for sections $D^{\prime},D^{\prime\prime},\tilde{D}$ of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$. Consequently, $(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}),\Delta,\epsilon)$ is a sheaf of $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras.
The Jet enveloping algebroid $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ as an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra: The notion of jet algebroid of a Lie algebroid $\mathcal{L}$ over $(X,\mathcal{O}_{X})$, is defined as the dual of the universal enveloping algebroid:
$\mathscr{J}(\mathcal{O}_{X},\mathcal{L}):=\mathscr{H}om_{\mathcal{O}_{X}}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}),\mathcal{O}_{X})$.
For each open set $U$ of $X$, we find the jet algebras $\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ of the $(\mathbb{K},\mathcal{O}_{X}(U))$-Lie-Rinehart algebra $\mathcal{L}(U)$ (see [7, 19, 2]). The jet algebroid $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ is the sheafification of the canonical presheaf:
$U\mapsto\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U)).$
One can dualize the structures on $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$. The product on $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ is induced from the coproduct of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ on each space of sections, which is locally defined by
$(\phi_{1}\phi_{2})(D):=\phi_{1}(D_{(1)})\phi_{2}(D_{(2)})$
for sections $\phi_{1},\phi_{2}\in\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ and a section $D\in\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$. By cocommutativity of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, this defines a commutative algebra structure on $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$. The unit for this multiplication is locally given by the left counit $\epsilon:\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathcal{O}_{X}$, since
$(\epsilon\phi)(D)=\epsilon(D_{(1)})\phi(D_{(2)})=\phi(\epsilon(D_{(1)})D_{(2)})=\phi(D)$
for a section $\phi$ in $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ and a section $D$ in $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$. The left and right $\mathcal{O}_{X}$-module structure on $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ provides $\mathcal{O}_{X}-\mathcal{O}_{X}$-bimodule structure on $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$.
The product on $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ descends to a sheaf homomorphism $m:\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\otimes_{\mathbb{K}_{X}}\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$. We can therefore dualize the product to obtain a coproduct $\Delta^{*}:\mathscr{J}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathscr{J}(\mathcal{O}_{X},\mathcal{L})\otimes_{\mathcal{O}_{X}}\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ locally defined as
$\phi(DD^{\prime})=:\Delta^{*}(\phi)(D\otimes D^{\prime})=\phi_{(1)}(D\phi_{(2)}(D^{\prime}))$
Associativity of the multiplication implies that $\Delta^{*}$ is coassociative. The counit for this coproduct is given by $\epsilon^{*}:\phi\mapsto\phi(1_{\mathscr{U}(\mathcal{O}_{X},\mathcal{L})})$. It follows that $(\mathscr{J}(\mathcal{O}_{X},\mathcal{L}),\Delta^{*},\epsilon^{*})$ is an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra.
Note 2.14.
The jet algebroid is a formal groupoid that serves as the formal exponentiation of the Lie algebroid [5]. In [18], the author studied this structure for the path algebroid $\mathcal{P}_{X}$ of a smooth variety $X$, referring to it as the formal path groupoid.
Note 2.15.
For an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra $(\mathcal{A},\Delta,\epsilon)$, the sheaf of premitive elements is determined by the following assignment (see [29])
$U\mapsto\{a\in\mathcal{A}(U)~{}|~{}\Delta(U)(a)=a\otimes 1+1\otimes a\},$
denoted by $\mathscr{P}(\mathcal{A})$, provides a sheaf of $(\mathbb{K}_{X},\mathcal{O}_{X})$-Lie-Rinehart algebras as follows:
$\mathfrak{a}:(\mathscr{P}(\mathcal{A}),[\cdot,\cdot]_{c})\rightarrow(\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X}),[\cdot,\cdot]_{c})$
is defined by $\mathfrak{a}(D)(f)=\epsilon(D~{}f)$ for sections $f\in\mathcal{O}_{X}$ and $D\in\mathscr{P}({\mathcal{A}})$.
Note 2.16.
For a smooth ((((holomorphic)))) Lie algebroid [1] i.e. for a locally free sheaf of Lie-Rinehart algebras of finite rank over a smooth (complex) manifold, the universal enveloping algebroid is a cocomplete locally graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra ((((of finite type)))) [29].
The tangent sheaf over the affine scheme $Spec(\mathbb{A}^{\mathbb{N}})$ ((((where $\mathbb{A}^{\mathbb{N}}:=\mathbb{K}[x_{i}]_{i\in\mathbb{N}})$ [9, 29] and the free Lie algebroid $\mathcal{P}_{X}$ over a smooth manifold $X$ [18] are locally free sheaf of Lie-Rinehart algebras ((((of infinite rank)))). Thus, their universal enveloping algebroids are cocomplete locally graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras ((((of infinite type)))) [29].
3. Lie algebroid Cohomology over $a$-spaces
Lie algebroids over algebraic spaces (shortly, $a$-spaces) are discussed in Section 2.1 by considering certain quasicoherent sheaves of Lie-Rinehart algebras over $a$-spaces $(X,\mathcal{O}_{X})$ (form an abelian category). The category of representations of a Lie algebroid has enough injectives (since the category of quasicoherent sheaves of left $\mathcal{O}_{X}$-modules is enough injectives [14]), useful for doing homological algebra.
The Chevalley-Eilenberg-de Rham cohomology (or, Lie algebroid cohomology) of a coherent Lie algebroid over a complex manifold or over a Noetherian separated scheme is expressed as derived functor $Ext$ [32, 3]. In [32], the result is stated for locally free Lie algebroid of finite rank over a complex manifold and in [3], the result is proved for locally free Lie algebroid of finite rank over a Noetherian separated scheme. Here, we construct a more general cochain complex of sheaves, which coincides with the cochain complexes of sheaves used in the classical contexts [21, 32, 1, 3]. For that we do not require the coherent (or locally finitely presented) condition on the underlying $\mathcal{O}_{X}$-module of a Lie algebroid. Notably, for analytic and algebraic geometric setups, the base space $X$ does not need to be non-singular. However, in that case, the tangent sheaf, which becomes a coherent Lie algebroid, does not retain the property of being a locally free Lie algebroid. For more details, see [32, 29]. Moreover, we present the result in a simplified form by using Čech cohomology. As an application of Lie algebroid cohomology we consider Chern classe of a module (locally free $\mathcal{O}_{X}$-module of finite rank) over the Lie algebroid.
To consider Lie algebroid cohomology for $\mathcal{L}$ over $(X,\mathcal{O}_{X})$ with coeffecient in some $\mathcal{O}_{X}$-module $\mathcal{E}$, we need to consider the followings.
3.1. Atiyah algebroid
For an (quasicoherent) $\mathcal{O}_{X}$-module $\mathcal{E}$, we form a Lie algebroid consisting of the sheaf of differential operators on $\mathcal{E}$ of order $\leq 1$ [18, 32, 3], defined by
$\mathcal{A}t(\mathcal{E})=\{D\in\mathscr{E}nd_{\mathbb{C}_{X}}(\mathcal{E})|D(fs)=fD(s)+\sigma_{D}(f)s~{}~{}\mbox{ for a unique}~{}~{}\sigma_{D}\in\mathcal{T}_{X},$
where $f\in\mathcal{O}_{X}$ and $s\in\mathcal{E}$}\}}, with the anchor map defined by
$\sigma:\mathcal{A}t(\mathcal{E})\rightarrow\mathcal{T}_{X}$ where $D\mapsto\sigma_{D}$
and the Lie bracket is commutator bracket. This Lie algebroid structure is so-called Atiyah algebroid of the $\mathcal{O}_{X}$-module $\mathcal{E}$. It provides a short exact sequence (s.e.s.) of Lie algebroids over $(X,\mathcal{O}_{X})$ (an abelian Lie algebroid extension) as
[…]
In particular, when $\mathcal{E}=\mathcal{O}_{X}$ for a non singular $a$-space $X$, we have $\mathcal{A}t(\mathcal{O}_{X})\cong\mathcal{O}_{X}\oplus\mathcal{T}_{X}$, equals to the sheaf differential operators of order $\leq 1$ with scalar symbol. The universal enveloping algebroid of the Atiyah algebroid $\mathcal{A}t(\mathcal{O}_{X})$ is the sheaf of twisted differential operators over $X$ [2].
A connection $\nabla:\mathcal{T}_{X}\rightarrow\mathcal{A}t(\mathcal{E})$ that satisfies $\sigma\circ\nabla=Id_{\mathcal{T}_{X}}$ provides a splitting of the s.e.s. (1) as $\mathcal{O}_{X}$-modules. If its curvature is zero then the s.e.s. splits as Lie algebroids. More generally, for a Lie algebroid $(\mathcal{L},[\cdot,\cdot],\mathfrak{a})$ on an $a$-space $(X,\mathcal{O}_{X})$, a $\mathcal{L}$-connection on $\mathcal{E}$ is defined by an $\mathcal{O}_{X}$-linear map (see [32, 2, 3])
(2) $\nabla:\mathcal{L}\rightarrow\mathcal{A}t(\mathcal{E})$
$\hskip 14.22636ptD\mapsto\nabla_{D}$
satisfying the Leibniz rule $\nabla_{D}(f~{}s)=f~{}\nabla_{D}(s)+\mathfrak{a}(D)(f)~{}s$ for sections $f\in\mathcal{O}_{X}$, $D\in\mathcal{L}$ and $s\in\mathcal{E}$ ((((to ensure compatibility, we choose $\sigma\circ\nabla=\mathfrak{a}$, i.e. $\sigma_{\nabla_{D}}=\mathfrak{a}(D)$ for $D\in\mathcal{L})$. Equivalently, it is described by a $\mathbb{K}_{X}$-linear map
$\nabla:\mathcal{E}\rightarrow\Omega_{\mathcal{L}}^{1}\otimes_{\mathcal{O}_{X}}\mathcal{E}$
satisfying the Leibniz rule $\nabla(f~{}s)=f~{}\nabla s+\breve{\mathfrak{a}}(df)\otimes s$, where $\Omega_{\mathcal{L}}^{1}:=\mathscr{H}om_{\mathcal{O}_{X}}(\mathcal{L},\mathcal{O}_{X})$ for a Lie algebroid $\mathcal{L}$ and $\Omega_{X}^{1}:=\Omega_{\mathcal{T}_{X}}^{1}$ together with $\breve{\mathfrak{a}}:\Omega_{X}^{1}\rightarrow\Omega_{\mathcal{L}}^{1}$ is the dual of the anchor map.
The $\mathcal{L}$-connection on $\mathcal{E}$ is said to be flat if the map (2) is a Lie algebroid homomorphism (i.e. the $\mathcal{L}$-curvature is zero), i.e. the map (2) additionally satisfies
(3) $\nabla_{[D,D^{\prime}]}=[\nabla_{D},\nabla_{D^{\prime}}]_{c}$
In that case, $(\mathcal{E},\nabla)$ is said to be a representation of $\mathcal{L}$ or call it a $\mathcal{L}$-module.
Next, we consider some special cases where $X$ is non-singular $a$-spaces :
$(i)$ For $\mathcal{L}=\mathcal{T}_{X}$ and $\mathcal{E}$ is a locally free $\mathcal{O}_{X}$-module of finite rank (or vector bundle over $X$), we have the standard covariant derivative as a $\mathcal{T}_{X}$-connection on $\mathcal{E}$.
$(ii)$ For $\mathcal{L}=\mathcal{T}_{X}(-logY)$, if a connection exists on an $\mathcal{O}_{X}$-module $\mathcal{E}$, it is called a logarithmic connection on $\mathcal{E}$. It forms a meromorphic connection with poles along the divisor $Y$ [4, 32].
If $(\mathcal{E},\nabla)$ is a $\mathcal{L}$-module, then the Lie algebroid morphism
$\nabla:\mathcal{L}\rightarrow\mathcal{A}t(\mathcal{E})$
extends to an $\mathcal{O}_{X}$-linear homomorphism of $\mathbb{K}_{X}$-algebras (using Remark 2.8)
~$\tilde{\nabla}:\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathscr{E}nd_{\mathbb{K}_{X}}(\mathcal{E}),$
making $\mathcal{E}$ into a left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-module. If $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module and $\mathcal{E}$ be a $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-module, then the restriction of the action of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ to $\mathcal{L}$ (by using the canonical embedding of $\mathcal{L}$ in $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ as described in Note 2.9), provides a $\mathcal{L}$-module structure on $\mathcal{E}$. In this case, the category of $\mathcal{L}$-modules and the category of left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-modules are equivalent. It helps in studying homological algebra with $\mathcal{L}$-modules ($\mathcal{L}$ is a sheaf of non-associative $\mathbb{K}$-algebras) by treating them as modules over the associative $\mathbb{K}_{X}$-algebra $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$.
Notation: Let $\mathcal{O}$ be a sheaf of associative $\mathbb{K}$-algebras over a topological space $X$ with $\mathcal{O}$-modules $\mathcal{C}^{i}$, for $i\geq 0$. Consider a cochain complex $(\mathcal{C}^{\bullet},d)$ where $\mathcal{C}^{\bullet}:=\oplus_{i}\mathcal{C}^{i}$ is a graded $\mathcal{O}$-module and $d:\mathcal{C}^{\bullet}\rightarrow\mathcal{C}^{\bullet+1}$ is the co-boundary map, i.e. $d$ is a $\mathbb{K}_{X}$-linear map of degree $1$ satisfying $d^{2}=0$. Similarly, for a chain complex, instead of a co-boundary map we have a boundary map $\partial:\mathcal{C}^{\bullet}\rightarrow\mathcal{C}^{\bullet-1}$, a $\mathbb{K}_{X}$-linear map of degree $-1$ satisfying $\partial^{2}=0$. To address the following topics of discussion consistently, we consider such (co)chain complexes where $\mathcal{O}$-linearity of the (co)boundary maps is not necessarily required.
For a $\mathcal{O}$-module $\mathcal{C}$, we can canonically form the following sheaves of graded vector spaces $\wedge^{\bullet}_{\mathcal{O}}$ $\mathcal{C}:=\oplus_{i}\wedge^{i}_{\mathcal{O}}\mathcal{C}$ and $\mathcal{C}^{\otimes^{\bullet}}:=\oplus_{i}$ $\mathcal{C}^{\otimes^{i}}$ (forms exterior algebra and tensor algebra with respect to the products $\wedge$ and $\otimes$ over $\mathcal{O}$ respectively).
Note 3.1.
For a $\mathcal{O}$-module $\mathcal{F}$ on $X$, there exists an injective resolution $\mathcal{F}\overset{\sim}{\rightarrow}\mathcal{C}^{\bullet}(\mathcal{F})$ ((((a quasi-isomorphism) of $\mathcal{O}$-modules, known as the flabby Godement resolution of $\mathcal{F}$. Thus, for a complex of $\mathcal{O}$-modules $\mathcal{F}^{\bullet}$, we consider the associated bicomplex of sheaves $\mathcal{C}^{\bullet}(\mathcal{F}^{\bullet})=(\mathcal{C}^{p}(\mathcal{F}^{q}))$ $(p,q\in\mathbb{N}\cup\{0\})$ of injective resolutions. The original complex is embedded in the total complex $\mathcal{K}^{\bullet}=tot(\mathcal{C}^{\bullet}(\mathcal{F}^{\bullet}))$, and this embedding is a quasi-isomorphism. The cohomology of the associated complex $\mathcal{K}^{\bullet}(X)=tot(\mathcal{C}^{\bullet}(\mathcal{F}^{\bullet}))(X)$ of global sections is called the hypercohomology of $\mathcal{F}^{\bullet}$, and denoted by $\mathbb{H}^{\bullet}(X,\mathcal{F}^{\bullet})$ (see [8]).
Note 3.2.
Denote the category of cochain complexes of sheaves of $\mathcal{O}_{X}$-modules on an $a$-space $(X,\mathcal{O}_{X})$ by $Shv(X)^{\bullet}$ and the category of $\mathbb{K}$-vector spaces by $Vect$.
For $\mathcal{F}^{\bullet}\in Shv(X)^{\bullet}$, the $k$-th cohomology sheaf of the complex $\mathcal{F}^{\bullet}$ is
$\mathscr{H}^{k}(\mathcal{F}^{\bullet}):=\mathscr{K}er(\mathcal{F}^{k}\rightarrow\mathcal{F}^{k+1})/{\mathscr{I}m(\mathcal{F}^{k-1}\rightarrow\mathcal{F}^{k})}$
((((where $\mathcal{F}^{-1}:=0)$ for $k\in\mathbb{N}\cup\{0\}$, considers as quotient of $\mathbb{K}_{X}$-vector spaces.
Moreover, a map of complexes $\mathcal{F}^{\bullet}\rightarrow\mathcal{G}^{\bullet}$ is a quasi-isomorphism if the induced map on cohomology sheaves $\mathscr{H}^{k}(\mathcal{F}^{\bullet})\rightarrow\mathscr{H}^{k}(\mathcal{G}^{\bullet})$ is an isomorphism for all $k$.
The $k$-th hypercohomology is a functor $\mathbb{H}^{k}(X,-):Shv(X)^{\bullet}\rightarrow Vect$ for $k\in\mathbb{N}\cup\{0\}$ (see [8, 36] for details) satisfying the following two conditions:
A quasi-isomorphism of complexes $f^{\bullet}:\mathcal{F}^{\bullet}\rightarrow\mathcal{G}^{\bullet}$ induces an isomorphism $\mathbb{H}^{k}(X,f^{\bullet})$, and if $\mathcal{I}^{\bullet}$ is a complex of injective sheaves then $\mathbb{H}^{k}(X,\mathcal{I}^{\bullet})=H^{k}(\Gamma(X,\mathcal{I}^{\bullet}))$.
Denote the graded vector space $\bigoplus\limits_{n=0}^{\infty}\mathbb{H}^{n}(X,\mathcal{F}^{\bullet})$ as $\mathbb{H}^{\bullet}(X,\mathcal{F}^{\bullet})$.
Remark 3.3.
We use sheaf-theoretic generalizations of the well-known derived functors $Ext$ and $Cotor$, by sheafifying the standard cochain complexes ((((see [34], [19])))) and subsequently considering the associated hypercohomologies.
3.2. Chevalley-Eilenberg-de Rham complex and Koszul-Rinehart resolution
For a Lie algebroid $\mathcal{L}$ over $(X,\mathcal{O}_{X})$ with a representation $(\mathcal{E},\nabla)$, consider the cochain complex (consists with $\mathcal{O}_{X}$-modules with a $\mathbb{K}_{X}$-linear degree $1$ map), a generalization of the well known Chevalley-Eilenberg-de Rham complex [21]
(4) $\Omega^{\bullet}_{\mathcal{L}}(\mathcal{E}):=(\mathscr{H}om_{\mathcal{O}_{X}}(\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L},\mathcal{E}),d).$
It is the sheafification of the presheaf of cochain complexes [34, 39]
$U\mapsto(Hom_{\mathcal{O}_{X}(U)}(\wedge^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U),\mathcal{E}(U)),d_{U})=:\Omega_{\mathcal{L}(U)}^{\bullet}(\mathcal{E}(U)),$
where the differential $d_{U}$ associated with an open subset $U$ of $X$ is given by
$\begin{split}d_{U}(\omega)(D_{1}\wedge\cdots\wedge D_{k+1})&=\sum^{k+1}_{i=1}(-1)^{i+1}~{}\nabla^{U}_{D_{i}}(\omega(D_{1}\wedge\cdots\wedge\hat{D_{i}}\wedge\cdots\wedge D_{k+1}))\\ &+\sum_{i<j}(-1)^{i+j}~{}\omega([D_{i},D_{j}]\wedge D_{1}\wedge\cdots\wedge\hat{D_{i}}\wedge\cdots\wedge\hat{D_{j}}\wedge\cdots\wedge D_{k+1}),\end{split}$
where $D_{1},\dots,D_{k+1}\in\mathcal{L}(U)$ and $\omega\in\Omega^{k}_{\mathcal{L}(U)}(\mathcal{E}(U))$, and $\nabla^{U}:=\nabla(U)$ (see the notion in (2), (3)). Notice that, the differential $d$ (i.e. $d^{2}=0)$ is a $\mathbb{K}_{X}$-linear map, but not an $\mathcal{O}_{X}$-module homomorphism ((((satisfies the Leibniz rule: $d(f~{}\omega)=df\wedge\omega+f~{}d\omega$, for $f\in\mathcal{O}_{X}(U),\omega\in\Omega^{k}_{\mathcal{L}(U)}(\mathcal{E}(U)))$.
If both $\mathcal{L}$ and $\mathcal{E}$ are locally free $\mathcal{O}_{X}$-modules of finite rank, then we get
(5) $\mathscr{H}om_{\mathcal{O}_{X}}(\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L},\mathcal{E})\cong\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}\otimes_{\mathcal{O}_{X}}\mathcal{E},$
which implies $\Omega^{\bullet}_{\mathcal{L}}(\mathcal{E})\cong\Omega^{\bullet}_{\mathcal{L}}\otimes_{\mathcal{O}_{X}}\mathcal{E}$ where $\Omega^{\bullet}_{\mathcal{L}}:=\Omega^{\bullet}_{\mathcal{L}}(\mathcal{O}_{X})$. This complex is equivalent to the standard complex associated with a smooth Lie algebroid [21] or a holomorphic Lie algebroid [32]. In particular, we get the Chevalley-Eilenberg complex when $\mathcal{L}$ is a Lie algebroid over a point; and the de Rham complex $\Omega^{\bullet}_{X}$ when $\mathcal{L}=\mathcal{T}_{X}$ and $\mathcal{E}=\mathcal{O}_{X}$ over a non-singular (or smooth) $a$-space $X$.
The associated hypercohomology of the cochain complex (4) of $\mathcal{O}_{X}$-modules is called the Lie algebroid cohomology of $\mathcal{L}$ with coefficient in $\mathcal{E}$ and denoted by $\mathbb{H}^{\bullet}(\mathcal{L},\mathcal{E})$.
Consider the chain complex of left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-modules (call it by Koszul-Rinehart complex)
(6) $\mathscr{K}_{\mathcal{O}_{X}}^{\bullet}\mathcal{L}:=(\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\otimes_{\mathcal{O}_{X}}\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L},\partial),$
it is the sheafification of the presheaf of cochain complexes [34, 39]
$U\mapsto(\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U))\otimes_{\mathcal{O}_{X}(U)}\wedge^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U),\partial_{U})=:K^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$
where the differential $\partial_{U}$ associated with an open subset $U$ of $X$ is given by
$\begin{split}\partial_{U}(\tilde{D}\otimes D_{1}\wedge\cdots\wedge D_{k})&=\sum^{k}_{i=1}(-1)^{i-1}\tilde{D}\bar{D_{i}}\otimes D_{1}\wedge\cdots\wedge\hat{D_{i}}\wedge\cdots D_{k}\\ &+\sum\limits_{i<j}(-1)^{i+j}\tilde{D}\otimes[D_{i},D_{j}]\wedge D_{1}\wedge\cdots\wedge\hat{D_{i}}\wedge\cdots\wedge\hat{D_{j}}\wedge\cdots\wedge D_{k},\end{split}$
where $\tilde{D}\bar{D_{i}}$ define by the canonical map $\iota_{U}:\mathcal{L}(U)\rightarrow\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U))$, for all $\tilde{D}\in\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ and $D_{1},\dots,D_{k}\in\mathcal{L}(U)$. Notice that the boundary (or chain) map $\partial$ is $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-linear.
When $\mathcal{L}$ is locally free as $\mathcal{O}_{X}$-module, it provides a locally free resolution of $\mathcal{O}_{X}$ as a left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-module (implies the stalk-wise exactness), known as the Koszul-Rinehart resolution. That is, it is given by the (augmented) chain complex of left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-modules
(7) $\mathscr{K}_{\mathcal{O}_{X}}^{\bullet}\mathcal{L}\rightarrow\mathcal{O}_{X},$
defined by the counit map (or augmentation map) $\epsilon:\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathcal{O}_{X}$ (is given as $\epsilon(\tilde{D})=\tilde{D}(1)$ for $\tilde{D}\in\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$), provides the quasi-isomorphism.
Note 3.4.
In the last section we consider analogous resolutions for $\mathcal{O}_{X}$ in the category of left $\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}$-comodules, where we replace the complex $\mathscr{K}_{\mathcal{O}_{X}}^{\bullet}\mathcal{L}$ with its dual complex, the dual Koszul-Rinehart complex $\widetilde{\mathscr{K}}_{\mathcal{O}_{X}}^{\bullet}\mathcal{L}$ ((((sheafifying the local counterpart as described in [19] accordingly)))).
Remark 3.5.
In [32, 1, 3], the Lie algebroid cohomology have considered for a special kind of Lie algebroid $\mathcal{L}$ (consists by coherent sheaf or locally free sheaf of finite rank) over a Noetherian separated scheme or a complex manifold $(X,\mathcal{O}_{X})$. The Lie algebroid cohomology of $\mathcal{L}$ with values in a $\mathcal{L}$-module $\mathcal{E}$ (a coherent $\mathcal{O}_{X}$-module) is the hypercohomology of the complex $(\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}\otimes_{\mathcal{O}_{X}}\mathcal{E},$ $d)$. This appears as a special case (5) of the complex described in (4).
Notations: From now on we use a special type of open cover $\{U_{x}\mid x\in X\}$ of $X$ for a locally free Lie algebroid $\mathcal{L}$ where the restrictions $\mathcal{L}|_{U_{x}}$ are free $\mathcal{O}_{X}|_{U_{x}}$-modules for every $U_{x}$. These open sets $U_{x}$’s are called special open sets.
We now describe Lie algebroid cohomology in terms of a derived functor.
Theorem 3.6.
Let $\mathcal{L}$ be a locally free Lie algebroid over $(X,\mathcal{O}_{X})$ and $\mathcal{E}$ a representation of $\mathcal{L}$. Then we get an isomorphism of graded vector spaces
$\mathbb{H}^{\bullet}(\mathcal{L},\mathcal{E})\cong Ext^{\bullet}_{\mathscr{U}(\mathcal{O}_{X},\mathcal{L})}(\mathcal{O}_{X},\mathcal{E})$.
Proof.
For every $x\in X$, the space of sections $\mathcal{L}(U_{x})$ is a $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebra and a free module over $\mathcal{O}_{X}(U_{x})$. Thus, the Koszul-Rinehart resolution (7) is a locally free resolution of the sheaf of local ring $\mathcal{O}_{X}$.
Here, we consider the two naturally associated presheaf of cochain complexes
$U\mapsto\Omega_{\mathcal{L}(U)}^{\bullet}(\mathcal{E}(U))$,
$U\mapsto Hom_{\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U))}(K^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U),\mathcal{E}(U))$
We have a canonical morphism between these two presheaves of cochain complexes as $\mathbb{K}_{X}$-modules, which is a local isomorphism or stalk-wise isomorphism, using results from [34] as local descriptions (applying on each special open neighborhoods $U_{x}$’s of $x\in X$). Thus, the associated sheaves of cochain complexes are isomorphic as $\mathbb{K}_{X}$-modules (using Lemma $3.2$ and Lemma $3.3$ from [29]), i.e.
$\Omega^{\bullet}_{\mathcal{L}}(\mathcal{E})\cong\mathscr{H}om_{\mathscr{U}(\mathcal{O}_{X},\mathcal{L})}(\mathscr{K}_{\mathcal{O}_{X}}^{\bullet}\mathcal{L},\mathcal{E}).$
Hence, the Chevalley-Eilenberg-de Rham complex of $\mathcal{L}$ with coefficient in $\mathcal{E}$ is isomorphic to dual of the Koszul-Rinehart resolution of $\mathcal{O}_{X}$ by $\mathcal{E}$ in the category of left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-modules.
Therefore, the associated cohomology sheaves (see Note 3.2) are isomorphic, i.e.
$\mathscr{H}^{k}(\Omega^{\bullet}_{\mathcal{L}},\mathcal{E})\cong\mathscr{E}xt^{k}_{\mathscr{U}(\mathcal{O}_{X},\mathcal{L})}(\mathcal{O}_{X},\mathcal{E})$, for every $k$.
Now, applying the hypercohomology functor $\mathbb{H}^{\bullet}(X,-)$ (see Note 3.2) we get the required result (since $\mathbb{H}^{\bullet}(\mathcal{L},\mathcal{E})=\mathbb{H}^{\bullet}(X,\Omega^{\bullet}_{\mathcal{L}}(\mathcal{E}))$ and $Ext^{\bullet}_{\mathscr{U}(\mathcal{O}_{X},\mathcal{L})}(\mathcal{O}_{X},\mathcal{E})=\mathbb{H}^{\bullet}(X,\mathscr{H}om_{\mathscr{U}(\mathcal{O}_{X},\mathcal{L})}(\mathscr{K}_{\mathcal{O}_{X}}^{\bullet}\mathcal{L},\mathcal{E}))$). ∎
Note 3.7.
One can find the proof for a locally free Lie algebroid $\mathcal{L}$ of finite rank over a Noetherian separated scheme $(X,\mathcal{O}_{X})$ discussed in [3], by using the ideas of [34] on the level of stalks $\mathcal{L}_{x}$ (since $\mathcal{L}_{x}$ is a $(\mathbb{K},\mathcal{O}_{X,x})$-Lie-Rinehart algebra, projective (in fact, free) module over $\mathcal{O}_{X,x}$ for all $x\in X)$.
Corollary 3.8.
Let $(X,\mathcal{O}_{X})$ be a complex manifold or smooth algebraic variety and $Y$ a free divisor in $X$. Therefore, the Lie algebroid $\mathcal{T}_{X}(-logY)$ of logarithmic derivation is locally free $\mathcal{O}_{X}$-module (of finite rank). Using Note 2.11, the logarithmic de Rham cohomology (hypercohomology of the complex (8) described in Remark 4.1 and Remark 4.2 for some special cases, as given in (9)) can be expressed as
$\mathbb{H}^{\bullet}(X,\Omega^{\bullet}_{X}(logY))\cong Ext^{\bullet}_{\mathcal{D}_{X}(-logY)}(\mathcal{O}_{X},\mathcal{O}_{X}).$
Corollary 3.9.
Consider the path algebroid $\mathcal{P}_{X}$ for a non-singular special $a$-space $(X,\mathcal{O}_{X})$ and its universal enveloping algebroid $\mathbb{D}_{X}$ (see Remark 2.12). Note that a module over the free Lie algebroid $\mathcal{P}_{X}$ can be viewed as a vector bundle with connection and a $\mathbb{D}_{X}$-module describe some system of PDE on the space of paths (see [18] for details).
The Theorem 3.6 for the Lie algebroid $\mathcal{P}_{X}$ provides the isomorphism
$\mathbb{H}^{\bullet}(\mathcal{P}_{X},\mathcal{E})\cong Ext^{\bullet}_{\mathbb{D}_{X}}(\mathcal{O}_{X},\mathcal{E}).$
Corollary 3.10.
Let $\mathcal{A}$ be a cocomplete locally graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra [29] of finite or infinite type (for examples consider Note 2.16). Then the Lie algebroid $\mathscr{P}(\mathcal{A})$ (sheaf of primitive elements, see Note 2.15) is a locally free $\mathcal{O}_{X}$-module (of finite or infinite rank accordingly). Thus, we have $\mathcal{A}\cong\mathscr{U}(\mathcal{O}_{X},\mathscr{P}(\mathcal{A}))$ [29]. Therefore, applying Theorem 3.6 for $\mathcal{L}=\mathscr{P}(\mathcal{A})$ provides the isomorphism
$\mathbb{H}^{\bullet}(\mathscr{P}(\mathcal{A}),\mathcal{E})\cong Ext^{\bullet}_{\mathcal{A}}(\mathcal{O}_{X},\mathcal{E}).$
Remark 3.11.
In complex geometry ((((algebraic geometry)))), for a cochain complex of coherent sheaves $\mathcal{F}^{\bullet}$ over an analytic space ((((algebraic variety)))), we compute the hypercohomology $\mathbb{H}^{\bullet}(X,\mathcal{F}^{\bullet})$ via Čech cohomology $\check{H}^{\bullet}(\mathcal{U},\mathcal{F}^{\bullet})$ associated with some good open cover $\mathcal{U}=\{U_{i}\}_{i}$ of $X$ ((((since $\mathbb{H}^{\bullet}(X,\mathcal{F}^{\bullet})\cong\check{H}^{\bullet}(\mathcal{U},\mathcal{F}^{\bullet})$, given by a canonical isomorphism). Specifically, we provide a good open cover by connected Stein spaces (affine varieties) and we use Leray’s theorem [8, 33] and Cartan-Serre’s vanishing theorem [11] for sheaf cohomology.
When we consider a locally free Lie algebroid $\mathcal{L}$ of finite rank over an analytic space (algebraic variety) $(X,\mathcal{O}_{X})$, we can compute Lie algebroid cohomology as derived functor using Čech cohomology.
In both cases we have $\mathbb{H}^{\bullet}(\mathcal{L},\mathcal{E})\cong\check{H}^{\bullet}(\mathcal{U},\Omega^{\bullet}_{\mathcal{L}}(\mathcal{E}))$. Since over Stein spaces ((((affine varieties)))) $U_{i}$, $\mathcal{L}(U_{i})$ is a projective $\mathcal{O}_{X}(U_{i})$-module [28, 38], thus we get an isomorphism
$H^{\bullet}(\mathcal{L}(U_{i}),\mathcal{E}(U_{i}))\cong Ext^{\bullet}_{\mathcal{U}(\mathcal{O}_{X}(U_{i}),\mathcal{L}(U_{i}))}(\mathcal{O}_{X}(U_{i}),\mathcal{E}(U_{i}))$.
Hence,
$\check{H}^{\bullet}(\mathcal{U},\Omega^{\bullet}_{\mathcal{L}}(\mathcal{E}))\cong\check{H}^{\bullet}(\mathcal{U},\mathscr{H}om_{\mathscr{U}(\mathcal{O}_{X},\mathcal{L})}(\mathscr{K}_{\mathcal{O}_{X}}^{\bullet}\mathcal{L},\mathcal{E})).$
Also, note that the whole ideas works for the classical case of Lie algebroids over smooth manifolds (i.e. for smooth Lie algebroids). In [1], the Lie algebroid ((((hyper))))cohomology of holomorphic ((((algebraic)))) Lie algebroids over a complex manifold ((((smooth scheme)))) has expressed by Čech cohomology using a good open cover consisting with connected Stein manifolds ((((affine schemes)))). However, the case associated with analytic spaces is not considered. Additionally, this result can be extended to quasicoherent $\mathcal{O}_{X}$-modules.
More generally, for a cochain complex of quasicoherent sheaves $\mathcal{F}^{\bullet}$ over a Noetherian separated scheme, by considering an affine open cover $\mathcal{U}=\{U_{i}\}_{i}$ of $X$, we get a canonical isomorphism $\mathbb{H}^{\bullet}(X,\mathcal{F}^{\bullet})\cong\check{H}^{\bullet}(\mathcal{U},\mathcal{F}^{\bullet})$ [14]. Thus, we get a similar result for a locally free Lie algebroid over a Noetherian separated scheme.
3.3. Chern classes:
Lie algebroid (or Lie-Rinehart algebra) connections, curvatures and the associated Chern classes (one of the Characteristic classes) has been studied in the context of differential geometry [10] and algebric geometry [23, 24, 25]. We review and extend some of the facts in our context as follows.
Classically, the first Chern class produces a bijection between the set of all isomorphism classes of complex line bundles on a manifold $X$ and the cohomology group $H^{2}(X,\mathbb{Z})$. More generally, for any Lie-Rinehart algebra $L$ which is projective as a left $R$-module there is a one-to-one correspondence between isomorphism classes of projective $R$-modules of finite rank and the Lie-Rinehart cohomology group $H^{2}(L,R)$.
Consider a $(\mathbb{K},R)$-Lie-Rinehart algebra $L$ which is projective as an $R$-module. Let $(E,\nabla)$ be a $L$-module, then we have the isomorphism of $\mathbb{K}$-vector spaces
$H^{2}(L,E)\cong Ext^{1}(L,E,\nabla),$
where $Ext^{1}(L,E,\nabla)$ is the set of equivalence classes of extensions of $L$ by the flat connection $(E,\nabla)$ [25]. This result can be viewed in the algebraic geometry set up (over the affine scheme $SpecR$). Analogously, for a locally free Lie algebroid $\mathcal{L}$ over an $a$-space $(X,\mathcal{O}_{X})$ with a $\mathcal{L}$-module $(\mathcal{E},\nabla)$ we get the isomorphism of $\mathbb{K}$-vector spaces
$\mathbb{H}^{2}(\mathcal{L},\mathcal{E})\cong Ext^{1}(\mathcal{L},\mathcal{E},\nabla).$
Note that the Chern class of a finitely generated projective $R$-module $E$ with a $L$-connection $\nabla$ on $E$, defined by the trace of the curvature $R_{\nabla}$ associated with the connection $\nabla$ [24, 25]. If $R$ is a regular $\mathbb{K}$-algebra (implies $L:=Der_{\mathbb{K}}(R)$ is a finitely generated projective $R$-module) then by considering the covariant connection on $E$ we get the first Chern class given by trace of the curvature
$c_{1}(E)=Trace(R_{\nabla})\in H^{2}(Der_{\mathbb{K}}(R),R).$
Analogous result holds for a locally free $\mathcal{O}_{X}$-module $\mathcal{E}$ of finite rank over the affine scheme $X=SpecR$. We compute analogous result in the general set up by sheafifying the local descriptions given in [23, 24].
For a complex manifold $X$, a locally free $\mathcal{O}_{X}$-module $\mathcal{E}$ of finite rank (equivalently a holomorphic vector bundle over $X$) always have covariant derivative as a flat $\mathcal{T}_{X}$-connection on $\mathcal{E}$. An arbitrary $\mathcal{T}_{X}$-connection on $\mathcal{E}$ is given by an $\mathcal{O}_{X}$-linear map
$\nabla:\mathcal{T}_{X}\rightarrow\mathcal{E}nd_{\mathbb{C}_{X}}(\mathcal{E})$
satisfying the Leibniz rule, and its curvature is the $\mathcal{O}_{X}$-linear map
$R_{\nabla}:\wedge^{2}_{\mathcal{O}_{X}}\mathcal{T}_{X}\rightarrow\mathcal{E}nd_{\mathbb{C}_{X}}(\mathcal{E})$
defined as
$R_{\nabla}(D\wedge D^{\prime})=[\nabla_{D},\nabla_{D^{\prime}}]_{c}-\nabla_{[D,D^{\prime}]_{c}}$
for any two sections $D,D^{\prime}$ of $\mathcal{T}_{X}$. For a $\mathcal{L}$-connection $\nabla$ on $\mathcal{E}$, one verifies that the curvature $R_{\nabla}\in\mathscr{H}om_{\mathcal{O}_{X}}(\wedge^{2}_{\mathcal{O}_{X}}\mathcal{L},\mathcal{E}nd_{\mathcal{O}_{X}}(\mathcal{E}))$, where $\mathcal{L}$ is a Lie algebroid.
Consider the induced $\mathcal{T}_{X}$-connection $ad~{}\nabla=[\nabla,\cdot]$ on $\mathcal{E}nd_{\mathcal{O}_{X}}(\mathcal{E})$, a flat connection. Thus by considering the Chevalley-Eilenberg-de Rham complex with coefficient in $\mathcal{E}nd_{\mathcal{O}_{X}}(\mathcal{E})$, we get the cohomology class $[R_{\nabla}]\in\mathbb{H}^{2}(X,\mathcal{E}nd_{\mathcal{O}_{X}}(\mathcal{E}))$ (since $d^{2}(R_{\nabla})=0)$. Hence, the Chern class of the $\mathcal{T}_{X}$-module $(\mathcal{E},\nabla)$ is
$c_{1}(\mathcal{E})=Trace(R_{\nabla})\in\mathbb{H}^{2}(X,\mathcal{O}_{X})$
where $Trace:\mathcal{E}nd(\mathcal{E})\rightarrow\mathcal{O}_{X}$ is the trace map.
The restriction of a $\mathcal{T}_{X}$-connection to the Lie algebroid $\mathcal{L}:=\mathcal{T}_{X}(-logY)$ ((((i.e. $\Omega_{\mathcal{L}}^{1}:=\Omega_{X}^{1}(logY)$) induces a logarithmic connection which is a meromorphic connection with poles along the divisor $Y$. Moreover, if $Y$ is a free divisor, then for a locally free $\mathcal{O}_{X}$-module $\mathcal{E}$ of finite rank with a $\mathcal{L}$-connection we get the first Chern class $c_{1}(\mathcal{E})\in\mathbb{H}^{2}(\mathcal{L},\mathcal{O}_{X})$. Note that a $\mathcal{L}$-connection might not exist in general. We can extend the notion of Chern classes for $\mathcal{T}_{X}$ or $\mathcal{T}_{X}(-logY)$ to arbitary Lie algebroids.
4. Logarithmic de Rham cohomology
We compute algebraic (analytic) de Rham cohomology groups of a family of non-singular varieties $Y_{t}$ for $t\in\mathbb{C}\setminus\{0\}$ and their associated singular variety $Y:=Y_{0}$ (which appears as principal divisors [29]), i.e. cohomology of the Lie algebroid $\mathcal{T}_{Y_{t}}$ for $t\in\mathbb{C}$. Then, we compute hypercohomology of the logarithmic de Rham complexes [27], which simplifies to the Lie algebroid cohomology for $\mathcal{T}_{X}(-logY)$ [32], as we work over some free divisors $Y\subset X$ [4, 32].
We use the notion $\langle\{s_{1},\dots,s_{n}\}\rangle$ for $R$-module generated by $s_{1},\dots,s_{n}\in E$, where $R$ is a $\mathbb{K}$-algebra and $E$ is an $R$-module (also extend to sheaf theoretic settings).
$(1)$ Rectangular Hyperbolas : We consider the affine space $X:=\mathbb{A}^{2}$ ($\mathbb{C}^{2}$ with the Zariski topology) with its coordinate ring $\tilde{R}=\mathcal{O}_{X}(X)=\mathbb{C}[x,y]$. The principal ideal $I=\langle xy-t\rangle\subset\tilde{R}$ provides rectangular hyperbolas $Y_{t}:=V(I)$ (vanishing set of $I$) for every $t$ in $\mathbb{C}\setminus\{0\}$. For $t\in\mathbb{C}\setminus\{0\}$, $Y_{t}$ is a non singular affine variety with its coordinate ring $R_{t}:=\mathcal{O}_{Y_{t}}(Y_{t})=\mathbb{C}[x,y]/\langle xy-t\rangle$.
Here, $Y:=Y_{1}$ is homeomorphic to $\mathbb{C}\setminus\{0\}$ and $\mathbb{C}\setminus\{0\}$ is same homotopy type of $S^{1}$. Thus, on considering the singular cohomology we get that
$H^{i}(Y,\mathbb{C})=H^{i}(S^{1},\mathbb{C})=\mathbb{C}$ for $i=0,1$
and $H^{i}(Y,\mathbb{C})=H^{i}(S^{1},\mathbb{C})=\{0\}$ for $i\geq 2$.
We can use algebraic de Rham theorem [8, 36] to get algebraic de Rham cohomology of $Y$ by the singular cohomology of $S^{1}$. Now we recall the computation which helps us to follow the associated singular case when $t=0$ (which is a normal crossing divisor).
[…]
$\Omega^{1}_{R}:=\frac{\langle\{dx,dy\}\rangle}{\langle df\rangle}=\langle dx\rangle.$
Note that $\frac{dx}{x}|_{Y}$ is an algebraic differential $1$-form on $Y$ (here $\Omega^{1}_{R}$ is a free $R$-module of rank $1$). The space of algebraic differential $2$-forms on $Y$ is $\Omega^{2}_{R}:=\wedge^{2}_{R}\Omega^{1}_{R}$. Now, $dx\wedge dy=dx\wedge-\frac{dx}{x^{2}}=0$ on $Y$ (since $y=\frac{1}{x}$ in $Y$) and thus we get $\Omega^{2}_{R}=\{0\}$.
Thus, the algebraic de Rham complex of $Y$ (or $SpecR$) is
$R\overset{d^{0}}{\longrightarrow}\Omega^{1}_{R}\overset{d^{1}}{\longrightarrow}0\cdots,$
where the first differential is given by $x\mapsto dx$ and $\frac{1}{x}\mapsto-\frac{dx}{x^{2}}$. Here, $\frac{dx}{x}$ is in the kernel of $d^{1}$ but $\frac{dx}{x}$ is not in the image of $d^{0}$ (the only possible choice of preimage is $logx$, but it is not a polynomial or regular function). It follows that $H^{1}_{dR}(Y)$ is a $\mathbb{C}$-vector space of dimension $1$. Now, $H_{dR}^{0}(Y)\cong\mathbb{C}$ and $H_{dR}^{n}(Y)=\{0\}$ for $n\geq 2$.
This is studied in [15], as a topological invariant for non singular spaces, but the analogous computation for the associated singular case is not derived there.
Note that the affine algebraic set $Y$ can be considered as an analytic space. Moreover, this space can be viewed as a principal divisor [29, 32].
We consider the associated logarithmic de Rham cohomology [4, 27] in the context of complex geometry.
Remark 4.1.
Here, $\mathcal{O}_{X}$ is the sheaf of holomorphic functions on $X:=\mathbb{C}^{2}$ and $(Y:=V(\mathcal{I}),\mathcal{O}_{Y}:=\mathcal{O}_{X}/{\mathcal{I}})$ is the analytic space associated with the principal ideal sheaf $\mathcal{I}=\langle xy-1\rangle$ of $\mathcal{O}_{X}$. The sheaf of logarithmic derivations for $Y$ in $X$ is $\mathcal{T}_{X}(-logY)=\langle\{x\partial_{x}-y\partial_{y},f~{}\nabla f\}\rangle$ where $\nabla f=y\partial_{x}+x\partial_{y}$ ((((since $f=xy-1)$. The sheaf of logarithmic $1$-forms for $Y$ in $X$ (i.e. meromorphic $1$-forms $\omega$ of $X$ with poles along the divisor $Y$ such that $f\omega\in\Omega^{1}_{X}$ and $df\wedge\omega\in\Omega^{2}_{X}$ [27]) is
$\Omega^{1}_{X}(logY)=\mathscr{H}om_{\mathcal{O}_{X}}(\mathcal{T}_{X}(-logY),\mathcal{O}_{X})=\langle\{dx,dy,\frac{df}{f}\}\rangle,$
where $df=ydx+xdy$. Note that, the relation $y~{}dx+x~{}dy-(xy-1)~{}\frac{df}{f}=0$ holds.
The sheaf of logarithmic differential $2$-forms on $X$ associated with $Y$ is
$\Omega^{2}_{X}(logY):=\wedge^{2}_{\mathcal{O}_{X}}\Omega^{1}_{X}(logY)=\langle\{dx\wedge\frac{df}{f},dy\wedge\frac{df}{f},dx\wedge dy\}\rangle.$
To determine hypercohomology [8, 36] of the logarithmic de Rham complex
$\Omega^{\bullet}_{X}(logY):$ $\mathcal{O}_{X}\overset{d^{0}}{\longrightarrow}\Omega^{1}_{X}(logY)\overset{d^{1}}{\longrightarrow}\wedge^{2}_{\mathcal{O}_{X}}\Omega^{1}_{X}(logY)\overset{d^{2}}{\longrightarrow}\cdots$,
we have $\mathscr{K}er$ $d^{0}=\mathbb{C}_{X}$; $\mathscr{K}er$ $d^{1}$ is the $\mathbb{C}_{X}$-vector space generated by
$\{dg~{}|~{}g\in\mathcal{O}_{X}\}\cup\{\frac{df}{f}\}$, i.e. $\mathscr{K}er$ $d^{1}=\mathscr{I}m$ $d^{0}\oplus\mathbb{C}_{X}\{\frac{df}{f}\}$ (here $d^{0}=d)$;
and $\mathscr{K}er$ $d^{2}$ is the $\mathbb{C}_{X}$-vector space generated by $\{dg\wedge\frac{df}{f}~{}|~{}g\in\mathcal{O}_{X}\}\cup\{dx\wedge dy\}=\Omega^{2}_{X}(logY)$. Therefore,
$\mathscr{K}er$ $d^{2}=\mathbb{C}_{X}\{d^{1}(g~{}\frac{df}{f})~{}|~{}g\in\mathcal{O}_{X}\}\oplus\mathbb{C}_{X}\{d^{1}(x~{}dy)\}=\mathscr{I}m$ $d^{1}$.
Therefore, the hyperchomology groups (taking the space of global sections of the complex, since $X$ is a Stein manifold) are given by the following
$\mathbb{H}^{0}(X,\Omega^{\bullet}_{X}(logY))=\mathbb{C},~{}\mathbb{H}^{1}(X,\Omega^{\bullet}_{X}(logY))=\mathbb{C},~{}\mathbb{H}^{n}(X,\Omega^{\bullet}_{X}(logY))=\{0\}~{}(n\geq 2).$
On $Y$ we have $df=ydx+xdy=0$ implies $\frac{dx}{x}+\frac{dy}{y}=0$ ($x\neq 0,y\neq 0$) and by integrating it we get $xy=t$ are the solutions which parametrized by $t\in\mathbb{C}\setminus\{0\}$.
Thus, we can view the normal crossing divisor $Y_{0}:=\{(x,y)\in\mathbb{C}^{2}\mid xy=0\}$ (a singular analytic space and a free divisor) as deformation of family of rectangular hyperbolas $Y_{t}$ (appears in [31] as deformation of a scheme).
$(2)$ Normal Crossing Divisor : Here we compute the algebraic de Rham cohomology of the normal crossing divisor $Y_{0}$ which is a singular affine algebraic set in $X$ with the space of global section of its structure sheaf
$R_{0}:=\mathcal{O}_{Y_{0}}(Y_{0})=\frac{\mathbb{C}[x,y]}{\langle xy\rangle}.$
The algebraic de Rham complex for $Y_{0}$ (or $Spec$ $R_{0})$ is
[…]
For $Y_{0}$ we have $d(xy)=0$. Thus, $xdy=-ydx$, which implies $xdx\wedge_{R}dy=0$ or $ydx\wedge_{R}dy=0$. But we cannot conclude that $dx\wedge_{R}dy=0$ on $Y_{0}$.
[…]
$\Omega^{1}_{R_{0}}=\frac{\langle\{dx,dy\}\rangle}{\langle ydx+xdy\rangle}$
[…]
$\Omega^{2}_{R_{0}}=\frac{\langle dx\wedge dy\rangle}{\langle\{xdx\wedge dy,ydx\wedge dy\}\rangle}.$
Since the only possible choice for a cohomology class in $H^{2}_{dR}(Y_{0})$ is the algebraic $2$-form $dx\wedge dy=d^{1}(xdy)$ is cohomologous with zero element, thus $H^{2}_{dR}(Y_{0})=\{0\}$.
Note that $Y_{0}$ is of same homotopy type to a point, thus singular cohomologies of $Y_{0}$ vanishes in all higher dimensions ($\geq 1$). Fortunately, for this singular variety the algebraic de Rham cohomology and the singular cohomology are same (see [12]).
This approach does not provide a topological invariant for all singular algebraic varieties, to resolve this problem a different approach is considered in [15].
We consider the associated cohomology in complex geometry context (see [29]).
Remark 4.2.
Here $\mathcal{O}_{X}$ is the sheaf of holomorphic functions on $X:=\mathbb{C}^{2}$ and $(Y_{0}:=V(\mathcal{I}),\mathcal{O}_{Y}:=\mathcal{O}_{X}/{\mathcal{I}})$ is the analytic space associated with the ideal sheaf $\mathcal{I}=\langle xy\rangle$. Here, the sheaf of logarithmic derivation is $\mathcal{T}_{X}(-logY_{0})=\langle\{x\partial_{x},y\partial_{y}\}\rangle$ and the sheaf of logarithmic $1$-forms $\Omega^{1}_{X}(logY_{0})=\langle\{\frac{dx}{x},\frac{dy}{y}\}\rangle$. Both are locally free $\mathcal{O}_{X}$-modules of rank $2$. The sheaf of logarithmic $2$-forms for the divisor $Y_{0}$ is $\Omega^{2}_{X}(logY_{0})=\langle\{\frac{dx}{x}\wedge\frac{dy}{y}\}\rangle$, a locally free $\mathcal{O}_{X}$-module of rank $1$.
Note that, the sheaf $\mathcal{T}_{X}(-logY_{0})$ has a canonical Lie algebroid structure, and $Y_{0}$ is a free divisor [32, 29]. Since, $log$ $x$ and $log$ $y$ both are not in $\mathcal{O}_{X}$, thus $\frac{dx}{x}$ and $\frac{dy}{y}$ are not in $\mathscr{I}m$ $d^{0}$ but both are in $\mathscr{K}er$ $d^{1}$. Similarly, $\frac{dx}{x}\wedge\frac{dy}{y}$ is in $\mathscr{K}er$ $d^{2}$ but not in $\mathscr{I}m$ $d^{1}$. Other choices for cocycles are appears as coboundaries.
Thus, the hypercohomology of the logarithmic de Rham complex
(8) $\Omega^{\bullet}_{X}(logY):\mathcal{O}_{X}\overset{d^{0}}{\longrightarrow}\Omega^{1}_{X}(logY_{0})\overset{d^{1}}{\longrightarrow}\wedge^{2}_{\mathcal{O}_{X}}\Omega^{1}_{X}(logY_{0})\overset{d^{2}}{\longrightarrow}\cdots$
is $\mathbb{H}^{0}(X,\Omega^{\bullet}_{X}(logY_{0}))=\mathbb{C}$, $\mathbb{H}^{1}(X,\Omega^{\bullet}_{X}(logY_{0}))=\mathbb{C}^{2}$ and $\mathbb{H}^{2}(X,\Omega^{\bullet}_{X}(logY_{0}))=\mathbb{C}$, all higher cohomologies are zero.
Now, consider the holomorphic Lie algebroid cohomology [1]
(9) $H^{n}_{holo}(\mathcal{T}_{X}(-logY_{0})):=\mathbb{H}^{n}(X,\Omega^{\bullet}_{X}(logY_{0}))\cong H^{n}_{dR}(X\setminus Y_{0},\mathbb{C})$
(using the standard Lemma of Atiyah-Hodge [36]), similar results appear in [4].
Therefore, the standard Lie algebroid cohomology of $\mathcal{T}_{X}(-logY_{0})$ followed by the isomorphism (9) provides the analytic (algebraic) de Rham cohomology of the torus $\mathbb{T}^{2}$ in $\mathbb{R}^{4}\cong\mathbb{C}^{2}$ (since $X\setminus Y_{0}=\{(x,y)\in\mathbb{C}^{2}\mid x\neq\bar{0}$ and $y\neq\bar{0}\}=(\mathbb{C}\setminus\{\bar{0}\})\times(\mathbb{C}\setminus\{\bar{0}\})$ is a complex manifold and homotopic to $S^{1}\times S^{1}=\mathbb{T}^{2})$. This agrees with the logarithmic de Rham cohomology for $\Omega^{\bullet}_{X}(logY_{0})$ (described in (8)) as computed.
Note 4.3.
In general for smooth algebraic variety and complex manifold, to compute its cohomology we need to consider sheaf of de Rham complex because the space of global sections not necessarily captures the whole information. In these cases we can use Čech cohomology [8, 33] by considering an affine open cover or Stein open cover ((((known as good open cover)))) [1, 36] accordingly.
5. Hochschild Cohomology for a Lie algebroid over $a$-spaces
In this section, we define Hochschild cohomology of an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra (see Section 2.3), using (Hopf-)Hochschild cohomology of a left bialgebroid [19] as local descriptions. In particular, we describe the cases associated with universal enveloping algebroid and jet algebroid of a locally free Lie algebroid over an $a$-space. The cohomology groups can be computed using suitable standard complexes. We prove a version of Hochschild-Kostant-Rosenberg (HKR) theorem and dual HKR theorem. It is done by following the local counterpart as can be found in [19], where the authors proved (dual) HKR theorem for (finitely generated) projective Lie-Rinehart algebras. Moreover, we present the HKR theorem for the sheaf of logarithmic differential operators and the sheaf of noncommuative differential operators.
5.1. Hochschild hypercohomology of an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra
Let $(\mathcal{A},\Delta,\epsilon)$ be an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra. Consider the presheaf of cochain complexes of $\mathcal{O}_{X}$-modules
$U\mapsto C^{\bullet}(\mathcal{A}(U)),$
where $C^{\bullet}(\mathcal{A}(U))=((\mathcal{A}(U))^{\otimes_{\mathcal{O}_{X}(U)}^{\bullet}},b)$ is the Hochschild cochain complex of the $\mathcal{O}_{X}(U)/{\mathbb{K}}$-bialgebra $\mathcal{A}(U)$ with the differential $b$ defined for any $a_{1},\dots,a_{n}\in\mathcal{A}(U)$ as follows:
$b(a_{1}\otimes\cdots\otimes a_{n})=1\otimes a_{1}\otimes\cdots\otimes a_{n}+\sum^{n}_{i=1}(-1)^{i}a_{1}\otimes\cdots\otimes\Delta(a_{i})\otimes\cdots\otimes a_{n}$ $+(-1)^{n+1}a_{1}\otimes\cdots\otimes a_{n}\otimes 1.$
The sheafification of the presheaf provides a cochain complexes of $\mathcal{O}_{X}$-modules, we call it as Hochschild cochain complex for $\mathcal{A}$ and denote it by $\mathscr{C}^{\bullet}(\mathcal{A})$. Its associated hypercohomology is called Hochschild hypercohomology of (the $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra) $\mathcal{A}$, and denoted by $\mathbb{H}H^{\bullet}(\mathcal{A})$. To compute it, it is often useful to consider an appropriate resolution. The canonical choice is given as follows.
Consider the presheaf of cobar complex of left $\mathcal{A}$-comodules
$U\mapsto Cob^{\bullet}({\mathcal{A}(U)})$
where $Cob^{\bullet}({\mathcal{A}(U)}):=((\mathcal{A}(U))^{\otimes_{\mathcal{O}_{X}(U)}^{\bullet+1}},b^{\prime})$ is the cobar complex for $\mathcal{O}_{X}(U)$ in the category of left $\mathcal{A}(U)$-comodules with the differential $b^{\prime}$ defined for any $a_{0},\dots,a_{n}\in\mathcal{A}(U)$ as follows:
$b^{\prime}(a_{0}\otimes a_{1}\otimes\cdots\otimes a_{n})=\sum^{n}_{i=0}(-1)^{i}a_{0}\otimes\cdots\otimes\Delta(a_{i})\otimes\cdots\otimes a_{n}+(-1)^{n+1}a_{0}\otimes\cdots\otimes a_{n}\otimes 1.$
The sheafification of the presheaf $U\mapsto Cob^{\bullet}(\mathcal{A}(U))$, denotes as $\mathscr{C}ob^{\bullet}(\mathcal{A})$. It provides a resolution of $\mathcal{O}_{X}$ by cofree left $\mathcal{A}$-comodules (a quasi-isomorphism) given by the left $\mathcal{O}_{X}$-module structure on $\mathcal{A}$ as
$\mathcal{O}_{X}\rightarrow\mathscr{C}ob^{\bullet}(\mathcal{A}),$
call it cobar resolution of $\mathcal{A}$. Then by applying cotensor product functor $\mathcal{O}_{X}\Box_{\mathcal{A}}-$ (sheafifying the cotensor functor described in [19]) on it and considering hypercohomology (i.e. applying hypercohomology $\mathbb{H}^{\bullet}(X,-)$ functor on the induced cochain complex $\mathcal{O}_{X}\Box_{\mathcal{A}}\mathscr{C}ob^{\bullet}(\mathcal{A})$ of $\mathbb{K}_{X}$-vector spaces), we get the following isomorphism (sheafifying the local descriptions of relationship among Hochschild cochain complex and cobar resolution [19] and considering their associated hypercohomologies)
(10) $\mathbb{H}H^{\bullet}(\mathcal{A})\cong\mathbb{H}^{\bullet}(X,\mathcal{O}_{X}\Box_{\mathcal{A}}\mathscr{C}ob^{\bullet}(\mathcal{A}))=Cotor^{\bullet}_{\mathcal{A}}(\mathcal{O}_{X},\mathcal{O}_{X}).$
Now, we describe Hochschild cohomology for some special $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras, namely for the universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ [29] and the jet algebroid $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ [29] of a (locally free) Lie algebroid $\mathcal{L}$ over a special $a$-space (as mentioned in Note 2.2) $(X,\mathcal{O}_{X})$.
5.2. Hochschild hypercohomology of universal enveloping algebroid
First we recall some of the essential ideas related for our next topic of discussion.
The sheaf of $\mathcal{L}$-poly-vector fields on $X$ is defined as $\mathcal{T}^{poly}_{\mathcal{L}}(\mathcal{O}_{X}):=\bigoplus_{i}\wedge^{i}\mathcal{L}$. This generalizes the space of multisections of the tangent bundle (i.e. multi-vector fields) descried for different $a$-spaces occurs in geometries. It has a canonical sheaf of Gerstenhaber algebra structure on $X$. For the case of a smooth manifold $X$, Kontsevich introduced the (sheaf of) poly-differential operators on $X$. It induces a subcomplex of the Hochschild cochain complex for $\mathcal{O}_{X}$. Its analogue in the context of Lie algebroid is the sheaf of $\mathcal{L}$-poly differential operators $\mathcal{D}^{poly}_{\mathcal{L}}(\mathcal{O}_{X}):=\bigoplus_{i}\mathscr{U}(\mathcal{O}_{X},\mathcal{L})^{\otimes_{\mathcal{O}_{X}}^{i}}$. It has a canonical sheaf of Gerstenhaber algebra structure on $X$. In particular, when $\mathcal{L}=\mathcal{T}_{X}$ we denote these sheaves by the standard notions $\mathcal{T}^{poly}_{X}$ and $\mathcal{D}^{poly}_{X}$ respectively. The so-called HKR map is a quasi-isomorphism between $\mathcal{T}^{poly}_{X}$ and $\mathcal{D}^{poly}_{X}$ (of differential graded $\mathbb{K}_{X}$-vector spaces). See [7] for more generalities.
Using ideas from proof of HKR theorem (and canonical PBW coalgebra isomorphism) for Lie-Rinehart algebras [19] we get the following results. To state these results first we need to recall some notations associated with a $(\mathbb{K},R)$-Lie-Rinehart algebra $L$ and an $R/{\mathbb{K}}$-bialgebra $A$. The symmetric algebra of $L$ over $R$ (using underlying $R$-module structure) is denoted by $S_{R}L$ and $S_{R}L^{*}$ is the symmetric algbra for the $R$-module $L^{*}:=Hom_{R}(L,R)$, the universal enveloping algebra of the $(\mathbb{K},R)$-Lie-Rinehart algebra $L$ is denoted by $\mathcal{U}(R,L)$ and its dual is the jet algebra $\mathcal{J}(R,L)$, the cobar resolution of $A$ (using underlying $R$-coalgebra structure) is denoted by $Cob^{\bullet}(A)$ and the associated cohomology (applying the functor $R$ $\Box_{A}$ using cotensor product) is Hochschild cohomology of the $R/{\mathbb{K}}$-bialgebra $A$ denoted by $HH^{\bullet}(A)$ (in our cases $A$ is either $S_{R}L$, $\mathcal{U}(R,L)$ or $S_{R}L^{*}$, $\mathcal{J}(R,L)$ with canonical $R/\mathbb{K}$-bialgebra structure [26, 7, 5, 19]).
We know that the sheaf of symmetric algebras $\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}$ can be viewed as universal enveloping algebroid of the Lie algebroid $\mathcal{L}$ if the $\mathcal{O}_{X}$-module $\mathcal{L}$ is equipped with zero bracket and zero anchor (on each space of sections) [29]. Recall that we have PBW presheaf homomorphism of presheaf of $\mathcal{O}_{X}$-algebas given as
$\theta:S_{\mathcal{O}_{X}}\mathcal{L}\rightarrow\mathcal{U}(\mathcal{O}_{X},\mathcal{L})$
$f\otimes D_{1}\otimes\cdots\otimes D_{k}\mapsto\frac{1}{k!}f\underset{\sigma\in S_{k}}{\sum}\bar{D}_{\sigma(1)}\cdots\bar{D}_{\sigma(k)}$
where $f$ is a section of $\mathcal{O}_{X}$, $D_{i}$ is a section of $\mathcal{L}$ and $\bar{D}_{i}$ is its associated image in $\mathcal{U}(\mathcal{O}_{X},\mathcal{L})$ $(i=1,2,\dots,k)$. Sheafification of this map gives the generalized PBW map between the associated sheaves (without considering the graded quotient) [29]. In addition, if $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module then we get following results.
Lemma 5.1.
If a Lie algebroid $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module, then the generalized PBW map (or symmetrization map)
$\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}\rightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$
is an isomorphism of $\mathcal{O}_{X}$-coalgebras.
Proof.
First we take the PBW map associated with space of sections of $\mathcal{L}$ over each special open subset $U_{x}$ in $X$. It provides an $\mathcal{O}_{X}(U_{x})$-coalgebra isomorphism (between the cocommutative $\mathcal{O}_{X}(U_{x})$-coalgebras)
$S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})\cong\mathcal{U}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x})).$
It induces stalkwise isomorphisms and by using the sheafification functor over the underlying presheaf homomorphism we get the generalized PBW map
$\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}\rightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$
as an isomorphism of $\mathcal{O}_{X}$-coalgebras. ∎
Lemma 5.2.
If a Lie algebroid $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module, then there is an isomorphism of graded $\mathbb{K}$-vector spaces associated with each special open set $U_{x}$,
$HH^{\bullet}(\mathcal{U}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x})))\cong\wedge_{\mathcal{O}_{X}(U_{x})}^{\bullet}\mathcal{L}(U_{x}).$
Proof.
The $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebra $\mathcal{L}(U_{x})$ is a free $\mathcal{O}_{X}(U_{x})$-module for each special open set $U_{x}$. As in[19], we consider the isomorphism of cochain complexes associated with each special open set $U_{x}$ given as $Cob^{\bullet}(\mathcal{U}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))\overset{\cong}{\rightarrow}Cob^{\bullet}(S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})).$ The isomorphism induces an isomorphism of graded vector spaces
$HH^{\bullet}(S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x}))\cong HH^{\bullet}(\mathcal{U}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x})))$
associated with every special open set $U_{x}$ for $x\in X$. Now for an open set $U\subset X$ we have the anti-symmetrization (or skew-symmetrization) map
$Alt_{U}:\wedge_{\mathcal{O}_{X}(U)}^{\bullet}\mathcal{L}(U)\rightarrow(S_{\mathcal{O}_{X}(U)}\mathcal{L}(U))^{\otimes^{\bullet}}$ given by
$D_{1}\wedge\cdots\wedge D_{n}\mapsto\frac{1}{n!}\sum_{\sigma\in S_{n}}(-1)^{\sigma}D_{\sigma(1)}\otimes\cdots\otimes D_{\sigma(n)}$,
as well as the map
$P_{U}:(S_{\mathcal{O}_{X}(U)}\mathcal{L}(U))^{\otimes^{\bullet}}\rightarrow\wedge_{\mathcal{O}_{X}(U)}^{\bullet}\mathcal{L}(U)$ defined as
$\tilde{D_{1}}\otimes\cdots\otimes\tilde{D_{n}}\mapsto Pr_{U}(\tilde{D_{1}})\wedge\cdots\wedge Pr_{U}(\tilde{D_{n}})$,
Here $Pr_{U}:S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)\rightarrow S^{1}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)=\mathcal{L}(U)$ is the projection map on the direct summand $S^{1}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)=\mathcal{L}(U)=\wedge^{1}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$ and $D_{i}\in\mathcal{L}(U)$, $\bar{D_{i}}\in S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$ $(i=1,\dots,n)$.
These morphisms defines cochain equivalence (or quasi isomorphism)
$Cob^{\bullet}(S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x}))\overset{\sim}{\rightarrow}(\wedge_{\mathcal{O}_{X}(U_{x})}^{\bullet}\mathcal{L}(U_{x}),0)$ on each special open set $U_{x}$ in $X$ and using that we get the isomorphism of graded vector spaces
$HH^{\bullet}(\mathcal{U}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x})))\cong\wedge_{\mathcal{O}_{X}(U_{x})}^{\bullet}\mathcal{L}(U_{x})$.
Note: We consider the presheaf of cobar resolutions as described in Section 5.1. Consider the cobar complex for $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, denote it by $\mathscr{C}ob^{\bullet}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}))$ and the associated hypercohomology is isomorphic to the Hochschild hypercohomology of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, i.e. $\mathbb{H}H^{\bullet}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}))$. Similar notions are applicable for the jet algebroid $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ which we describe in the next section 5.3.
On each open set $U$ of $X$, we can define the anti-symmetrization map
$\widetilde{Alt}_{U}:\wedge^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)\rightarrow(\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U)))^{\otimes^{\bullet}}$
as the earlier case and we can check it is compatible with restrictions and differentials. Applying sheafification functor, we get homomorphism between associated cochain complex of $\mathcal{O}_{X}$-modules, we call it anti-symmetrization map (it depends on context). We use these results in the next theorem.
Lemma 5.3.
If $\mathcal{F}$ is a sheaf of $\mathcal{O}_{X}$-modules, then by considering the exterior algebra as a chain complex $\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{F}:=(\oplus_{i}\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F},0)$ we get a canonical isomorphism of $\mathbb{K}$-vector spaces
$\mathbb{H}^{k}(X,\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{F})\cong\underset{i+j=k}{\oplus}H^{j}(X,\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F})$
for every $k\geq 0$ (or $k$ is any non negative integer).
Proof.
Here we apply the notion of hypercohomology (Note 3.2) for the cochain complex of sheaves $\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{F}$. For each sheaf of ${\mathcal{O}_{X}}$-modules (or $\mathcal{O}_{X}$-module) $\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F}$, we have an injective resolution (given by the flabby Godement resolution) $\mathscr{C}^{\bullet}(\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F})$, provides a quasi-isomorphism
$\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F}\overset{\sim}{\rightarrow}\mathscr{C}^{\bullet}(\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F}).$
Thus, the sheaf cohomology $H^{\bullet}(X,\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F})$ of $\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F}$ ($i\in\mathbb{N}\cup\{0\})$ is isomorphic to the cohomology (in usual sense) of the complex of $\mathcal{O}_{X}(X)$-modules $\mathscr{C}^{\bullet}(\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F})(X)$.
The corresponding bicomplex for the complex of sheaves $\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{F}$ is $\mathscr{C}^{\bullet}(\mathcal{F}):=(\mathscr{C}^{j}(\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F}))_{i,j\geq 0}$. The original complex can be embedded in the total complex $\mathscr{K}^{\bullet}:=tot(\mathscr{C}^{\bullet}(\mathcal{F}))$. Moreover, this embedding is a quasi-isomorphism. The cohomology of the associated complex of global sections $\mathscr{K}^{\bullet}(X)=tot(\mathscr{C}^{\bullet}(\mathcal{F}))(X)$ is the hypercohomology of the complex $\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{F}$. Since here differential of the complex is zero, thus the computation of cohomology of the total complex is become a simpler one. It is directly expressed through the sheaf cohomologies $H^{j}(X,\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F})$ as stated in the Lemma.
Notations: We denote cotensor product with $\mathcal{O}_{X}(U)$ in the category of left $S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$-comodules as $\mathcal{O}_{X}(U)$ $\Box_{S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)}$- and the associated cohomology groups is given by the Cotor groups (functor) as $Cotor^{\bullet}_{S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)}(\mathcal{O}_{X}(U),-)$, for an open set $U$ of $X$. Instead of $S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$ we use $\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ when it requires.
Theorem 5.4.
((((Generalized HKR theorem)))) Let $\mathcal{L}$ be a locally free Lie algebroid over $(X,\mathcal{O}_{X})$. Then the anti-symmetrization map (known as HKR morphism)
$\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}\rightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})^{\otimes^{\bullet}}$
provides a quasi-isomorphism of the associated cochain complexes of sheaves. It induces an isomorphism of graded vector spaces
$\mathbb{H}H^{\bullet}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}))\cong\mathbb{H}^{\bullet}(X,\wedge^{\bullet}_{\mathcal{O}_{X}}{\mathcal{L}})$
Proof.
Assume first $\mathcal{L}$ to be a locally free Lie algebroid of finite rank. Since for each $x\in X$ we have an open set $U_{x}$ containing $x$ with $\mathcal{L}|_{U_{x}}$ is a free $\mathcal{O}_{X}|_{U_{x}}$-module, thus on each $U_{x}$ we use results from the proof of the HKR Theorem [19] for the $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebra $\mathcal{L}(U_{x})$. Then, for each special open sets $U_{x}$ we have isomorphisms of cochain complexes of left $S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})$-comodules
$(\mathcal{O}_{X}(U_{x})$ $\Box_{S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})}$ $\tilde{K}^{\bullet}_{U_{x}}\mathcal{L},$ $id_{\mathcal{O}_{X}(U_{x})}$ $\otimes_{\mathcal{O}_{X}(U_{x})}$ $\partial_{U_{x}})\overset{\phi_{U_{x}}}{\rightarrow}(\wedge^{\bullet}_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x}),0)$
compatible with restrictions, where $\tilde{K}^{\bullet}_{U}\mathcal{L}:=S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)\otimes_{\mathcal{O}_{X}(U)}\wedge^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$ and the differential $\partial_{U}:\tilde{K}^{\bullet}_{U}\mathcal{L}\rightarrow\tilde{K}^{\bullet+1}_{U}\mathcal{L}$ is described in [19], for any open set $U$ in $X$. Indeed, the dual Koszul-Rinehart complex for $\mathcal{L}(U)$ is given as $\tilde{K}^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U):=(\tilde{K}^{\bullet}_{U}\mathcal{L},$ $\partial_{U})$.
Now, an open set $U$ is expressed by taking a partition as $\cup_{x\in X}(U\cap U_{x})$ and thus we have cochain homomorphisms of left $S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$-comodules
$(\mathcal{O}_{X}(U)$ $\Box_{S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)}$ $\tilde{K}^{\bullet}_{U}\mathcal{L},$ $id_{\mathcal{O}_{X}(U)}$ $\otimes_{\mathcal{O}_{X}(U)}$ $\partial_{U})\overset{\phi_{U}}{\rightarrow}(\wedge^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U),0)$
compatible with restrictions. Then it induces a homomorphism on the associated presheaves of cochain complexes, provides stalkwise isomorphism. Thus, applying sheafification functor we get a canonical isomorphism of cochain complexes of left $\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}$-comodules
(11) $\mathcal{O}_{X}\Box_{\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}}\widetilde{\mathscr{K}}^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}=(\mathcal{O}_{X}\Box_{\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}}(\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}\otimes\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}),id_{\mathcal{O}_{X}}\otimes_{\mathcal{O}_{X}}\partial)\overset{\phi}{\rightarrow}(\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L},0).$
To compute Hochschild hypercohomology $\mathbb{H}H^{\bullet}(\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L})$ via a derived functor, we consider the standard cobar complex associated with each open set $U$, denoted as $Cob^{\bullet}(S_{\mathcal{O}_{X}(U)}\mathcal{L}(U))$ (considering $\mathcal{A}(U)=S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$ as described in Section 5.1 ).
We have two presheaf of vector spaces, one is presheaf of cobar complexes of the sheaf $\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}$ of $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra, defined as
$U\mapsto Cob^{\bullet}(S_{\mathcal{O}_{X}(U)}\mathcal{L}(U))$
with its restriction morphism induced from the restrictions of the sheaf $\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}$, and another one is the presheaf of Koszul-Rinehart complex of $\mathcal{O}_{X}$ in the category of left $S_{\mathcal{O}_{X}}\mathcal{L}$-comodules, defined as
$U\mapsto\widetilde{K}_{\mathcal{O}_{X}(U)}^{\bullet}\mathcal{L}(U).$
These two presheaves are quasi isomorphic on special open sets $U_{x}$ for every $x\in X$ (for the $C^{\infty}$ case these holds for every open set $U$). In stalkwise, it induces quasi-isomorphism and by applying sheafification functor we get quasi-isomorphism (chain homotopy)
(12) $\mathscr{C}ob^{\bullet}(\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L})\overset{\sim}{\rightarrow}\widetilde{\mathscr{K}}_{\mathcal{O}_{X}}^{\bullet}\mathcal{L}$
(using the alternating map and projection map (5.2), we have quasi isomorphism
$S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})\otimes_{\mathcal{O}_{X}(U_{x})}\wedge^{\bullet}_{\mathcal{O}_{X}(U_{x})}(\mathcal{L}(U_{x}))\overset{id\otimes Alt_{U_{x}}}{\longrightarrow}S(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))^{\otimes^{(\bullet+1)}}).$
Then on taking cotensor product by $\mathcal{O}_{X}$ from left (as standard left $\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}$-comodule) to the cochain complexes (12) of sheaves of left $\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}$-comodules we get from (11)
$\mathcal{O}_{X}\Box_{\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}}$ $\mathscr{C}ob^{\bullet}(\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L})\overset{\sim}{\rightarrow}(\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L},0)$.
Applying hypercohomology functor $\mathbb{H}^{\bullet}(X,-)$, we get Hochschild hypercohomology of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ (using Lemma 5.1) through derived functor Cotor as
$\mathbb{H}H^{\bullet}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}))\cong Cotor^{\bullet}_{\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}}(\mathcal{O}_{X},\mathcal{O}_{X})\overset{\cong}{\rightarrow}\mathbb{H}^{\bullet}(X,\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L})$.
An alternative approach for proof of the theorem for finite rank case is given as follows.
Consider the presheaves of graded vector spaces
$U\mapsto HH^{\bullet}(S_{\mathcal{O}_{X}(U)}\mathcal{L}(U))$,
$U\mapsto\wedge_{\mathcal{O}_{X}(U)}^{\bullet}$ $\mathcal{L}(U)$
are isomorphic on each special open sets $U_{x}$’s (see Lemma 5.2). The associated presheaf homomorphism is induced from the homomorphisms $Alt_{U}$ and $P_{U}$ (described in the proof of the Lemma 5.2). Thus, by considering sheafifications (of these two presheaf of cohomology spaces which are isomorphic stalkwise) we get an isomorphism in the associated cohomology sheaves (described in Definition 3.2) as
$\mathscr{H}H^{\bullet}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}))\cong\mathscr{H}H^{\bullet}(\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L})\cong\wedge_{\mathcal{O}_{X}}^{\bullet}\mathcal{L}.$
Next on the induced hypercohomology groups (using Note 3.2) we have
$\mathbb{H}H^{\bullet}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}))\cong\mathbb{H}H^{\bullet}(\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L})\cong\mathbb{H}^{\bullet}(X,\wedge_{\mathcal{O}_{X}}^{\bullet}\mathcal{L})$.
In general case, where $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module of infinite rank (quasicoherent sheaf), there exist a filtered ordered set $J$ as well as an inductive system of free $\mathcal{O}_{X}(U_{x})$-modules $\{{\mathcal{L}(U_{x})}_{j}\mid j\in J\}$ such that
[…]
for each special open set $U_{x}$ (using results from [19]). Since both $\mathbb{H}H^{\bullet}$ (which is isomorphic to the derived functor $Cotor^{\bullet}$) and the functor $\mathcal{S}$ commute with inductive limits over a filtered ordered set, we get the result using the result follows from the locally free $\mathcal{O}_{X}$-module of finite rank case (local counterpart is described in [19]). ∎
Remark 5.5.
The HKR morphism is not a ring isomorphism, but composing the HKR-morphism together with the Todd genus provides a ring isomorphism (moreover it provides a canonical Gerstenhaber algebra isomorphism). It is required in the study of formality (or quantization) for Lie algebroids [7].
Corollary 5.6.
By applying the Lemma 5.3, the above HKR theorem reduces to
$\mathbb{H}H^{\bullet}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}))\cong\underset{i,j}{\oplus}H^{j}(X,\wedge^{i}_{\mathcal{O}_{X}}\mathcal{L}).$
Corollary 5.7.
In the special case of $\mathcal{L}=\mathcal{T}_{X}$ when $X$ is non-singular (or smooth), we get isomorphism of graded vector spaces
$\mathbb{H}H^{\bullet}(\mathcal{D}_{X})=\mathbb{H}^{\bullet}(X,\mathcal{D}^{poly}_{X})\cong\mathbb{H}^{\bullet}(X,\mathcal{T}^{poly}_{X})$.
Corollary 5.8.
Using Note 2.11 for a free divisor $Y$ in $X$, we get isomorphism of graded vector spaces
$\mathbb{H}H^{\bullet}(\mathcal{D}_{X}(-logY))\cong\mathbb{H}^{\bullet}(X,\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{T}_{X}(-logY)).$
Corollary 5.9.
((((Noncommutative analogue of HKR Theorem)))) The free Lie algebroid $\mathcal{P}_{X}$ can be identified with sheaf of regular noncommutative vector fields on $X$ and we have seen already that the sections of its universal enveloping algebra $\mathbb{D}_{X}$ are noncommutative differential operators on $X$ ((((see Remark 2.12)))). Thus, in this case from the HKR theorem we get a canonical isomorphism
$\mathbb{H}H^{\bullet}(\mathbb{D}_{X})\cong\mathbb{H}^{\bullet}(X,\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{P}_{X})$
Corollary 5.10.
For a cocomplete graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra $\mathcal{A}$ ((((of finite or infinite type)))), the isomorphism $\mathcal{A}\cong\mathscr{U}(\mathcal{O}_{X},\mathscr{P}(\mathcal{A}))$ holds [29] and thus we get the graded vector space isomorphism
$\mathbb{H}H^{\bullet}(\mathcal{A})\cong\mathbb{H}^{\bullet}(X,\wedge^{\bullet}_{{\mathcal{O}_{X}}}\mathscr{P}(\mathcal{A})).$
5.3. Hochschild hypercohomology of jet algebroid
Here, we present a dual version of the HKR theorem by considering Hochschild cohomology of a jet algebroid $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ for a Lie algebroid $\mathcal{L}$ over $(X,\mathcal{O}_{X})$ (see the part (2.13) mentioned in Section 2).
Note 5.11.
There is a canonical left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-module structure on $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ constructed as follows ((((see [7, 19, 6] for the local descriptions)))):
A canonical flat $\mathcal{L}$-connection, called Grothendieck connection is given by
(13) $\nabla_{{D}}(\phi)(D^{\prime}):=\tilde{D}(\phi(D^{\prime}))-\phi(\bar{D}~{}D^{\prime})$
for all sections ${D}\in\mathcal{L}$, $\phi\in\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ and $D^{\prime}\in\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, where $\tilde{D}:=\mathfrak{a}(D)$ and $\bar{D}:=\iota_{\mathcal{L}}(D)$. Thus, we get a $\mathcal{L}$-module structure on $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$. We can extend it to the induced $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-module structure on $\mathcal{J}(R,L)$.
Theorem 5.12.
((((Generalized dual HKR theorem)))) Let $\mathcal{L}$ be a locally free Lie algebroid over $(X,\mathcal{O}_{X})$ which is of finite rank. Then the anti-symmetrization map ((((dual HKR morphism))))
$\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L^{*}}\rightarrow\mathscr{J}(\mathcal{O}_{X},\mathcal{L})^{\otimes^{\bullet}}$
provides a quasi-isomorphism between the associated cochain complexes of sheaves. It induces canonical isomorphism of graded vector spaces
$\mathbb{H}H^{\bullet}(\mathscr{J}(\mathcal{O}_{X},\mathcal{L}))\cong\mathbb{H}^{\bullet}(X,\Omega^{\bullet}_{\mathcal{L}}).$
Proof.
For a locally free Lie algebroid $\mathcal{L}$ over $(X,\mathcal{O}_{X})$ of finite rank (say $r$, for some $r\in\mathbb{N}$), the dual of $\mathcal{L}$, denoted by $\mathcal{L^{*}}$, is a locally free $\mathcal{O}_{X}$-module of the same rank. Thus, $\mathcal{L^{*}}(U_{x})$ has a basis $\{w_{1},\dots,w_{r}\}$ (say) for each special open set $U_{x}$ around $x\in X$. Since $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ is a commutative $\mathcal{O}_{X}$-algebra, thus $\mathcal{J}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))\cong\mathcal{O}_{X}(U_{x})[[w_{1},\dots,w_{r}]]$ as $\mathcal{O}_{X}(U_{x})$-algebra. Hence, by applying sheafification functor we get the isomorphism of $\mathcal{O}_{X}$-algebras as $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})\cong\widehat{\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L^{*}}}$ (this sheaf of symmetric algebras is formally completed with respect to the degree) [5, 2]. Now, consider the dual Koszul-Rinehart resolution of $\mathcal{O}_{X}$ in the category of left $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$-comodules (by using local counterpart from [19]) is
$\mathcal{O}_{X}\hookrightarrow(\mathscr{J}(\mathcal{O}_{X},\mathcal{L})\otimes_{\mathcal{O}_{X}}\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L^{*}},\nabla)=:\bar{\mathscr{K}}^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*},$
where $\nabla$ is the Grothendieck connections (Note 5.11), sheafification of the presheaf of canonical left $\mathcal{L}(U)$-connection $U\mapsto\nabla_{U}$ on $\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))$. It is basically the sheafification of the presheaves of cochain complex
$U\mapsto\{\mathcal{O}_{X}(U)\hookrightarrow\bar{K}^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)^{*}\},$
where $\bar{K}^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)^{*}=\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))\otimes_{\mathcal{O}_{X}(U)}\wedge^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)^{*}$. The unit of $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ as an $\mathcal{O}_{X}$-algebra provides the morphism from $\mathcal{O}_{X}$ to $\bar{\mathscr{K}}^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}$.
Applying cotensor product by $\mathcal{O}_{X}$ (with canonical $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$-comodule structure, which induces from the $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-module structure on $\mathcal{O}_{X}$) to the dual Koszul-Rinehart complex $\bar{\mathscr{K}}^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}$, we get the canonical isomorphism (sheafifying local descriptions from [19]) of cochain complexes of sheaves of graded vector spaces as
(14) $\mathcal{O}_{X}\Box_{\mathscr{J}(\mathcal{O}_{X},\mathcal{L})}\bar{\mathscr{K}}^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}\cong(\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L^{*}},d).$
To describe the isomorphism (14) in an explicit way, we use the following steps.
$\mathcal{O}_{X}(U)\Box_{\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))}(\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))\otimes_{\mathcal{O}_{X}(U)}\wedge^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)^{*})\cong\wedge^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)^{*}.$
Since the unit $1_{U}\in\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ is given by the counit $\epsilon_{U}$ of $\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U))$, the induced differential is exactly the Lie-Rinehart coboundary $d_{U}$ (4), on an open set $U\subset X$. These isomorphisms are compatible with restrictions.
Now consider the canonical presheaves from the above complexes associated with each open sets, which are isomorphic on each special open sets and compatible with the natural restrictions. Then considering there sheafifications and using the above isomorphisms, we get isomorphism between the complexes of sheaves, described in (14).
Since $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module of finite rank, there exists open sets $U_{x}$ around each point $x\in X$, where $\mathcal{L}|_{U_{x}}$ is free $\mathcal{O}_{X}|_{U_{x}}$-module of finite rank. Thus, for each $U_{x}$ we can use results from the proof of the dual HKR Theorem [19] for the $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebra $\mathcal{L}(U_{x})$. Here, we get
$\wedge^{\bullet}_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})^{*}\cong(\wedge^{\bullet}_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x}))^{*}.$
Hence, $\mathscr{H}om_{\mathcal{O}_{X}}(\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L},\mathcal{O}_{X})\cong\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L^{*}}$ and the Chevalley-Eilenberg-de Rham complex of $\mathcal{L}$ is $\Omega^{\bullet}_{\mathcal{L}}\cong(\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L^{*}},d)$. Therefore, we get (using isomorphism (14))
(15) $\mathcal{O}_{X}\Box_{\mathscr{J}(\mathcal{O}_{X},\mathcal{L})}\bar{\mathscr{K}}^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}\cong\Omega^{\bullet}_{\mathcal{L}}.$
To express Hochschild hypercohomology of $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ as derived functor, we need to consider standard cobar resolution of $\mathcal{O}_{X}$ as $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$-comodules and cotensor it with $\mathcal{O}_{X}$ (putting $\mathcal{A}=\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ in the relation appears in (10)). Thus,
(16) $\mathbb{H}H^{\bullet}(\mathscr{J}(\mathcal{O}_{X},\mathcal{L}))\cong Cotor^{\bullet}_{\mathscr{J}(\mathcal{O}_{X},\mathcal{L})}(\mathcal{O}_{X},\mathcal{O}_{X}).$
Since for each open set $U_{x}$ of $X$, the $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebras $\mathcal{L}(U_{x})$ is finitely generated projective (in fact free), thus we get quasi-isomorphisms
$\mathcal{J}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))^{\otimes(\bullet+1)}\overset{id_{U_{x}}\otimes P_{U_{x}}}{\longrightarrow}\mathcal{J}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))\otimes_{\mathcal{O}_{X}(U_{x})}\wedge^{\bullet}_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})^{*},$
$\mathcal{J}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))\otimes_{\mathcal{O}_{X}(U_{x})}\wedge^{\bullet}_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})^{*}\overset{id_{U_{x}}\otimes Alt_{U_{x}}}{\longrightarrow}(\mathcal{J}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x})))^{\otimes(\bullet+1)},$
for each $U_{x}$ associated with $x\in X$, where $P_{U_{x}}$ is given by the cannonical projections $pr_{1}:\mathcal{J}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))\cong\widehat{S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})^{*}}\rightarrow\mathcal{L}(U_{x})^{*}$ and the anti-symmetrization map $Alt_{U_{x}}:\wedge^{\bullet}_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})^{*}\rightarrow\mathcal{J}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))^{\otimes^{\bullet}}$ for every special open set $U_{x}$ of $X$ (for smooth manifold we get these quasi isomorphisms for each open sets). Sheafification of the associated presheaves provides the quasi-isomorphism between the cobar complex $Cob^{\bullet}(\mathscr{J}(\mathcal{O}_{X},\mathcal{L}))$ and the dual Koszul-Rinehart complex $\bar{\mathscr{K}}^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}$.
Applying hypercohomology functor $\mathbb{H}^{\bullet}(X,-)$ we get (from the isomorphism (15))
(17) $Cotor^{\bullet}_{\mathscr{J}(\mathcal{O}_{X},\mathcal{L})}(\mathcal{O}_{X},\mathcal{O}_{X})\cong\mathbb{H}^{\bullet}(X,\Omega^{\bullet}_{\mathcal{L}}).$
Thus, using the isomorphisms (16) and (17), the Hochschild hypercohomology groups of jet algebroid $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ is expressed by the Chevally-Eilenberg-de Rham hypercohomology of $\mathcal{L}$. ∎
Remark 5.13.
In [5, 6], a dual version of the (twisted) HKR theorem for Lie algebroids [7] (with rich algebraic structures) is considered in-order to study a notion of precalculus upto homotopy in the Lie algebroids context.
Corollary 5.14.
In particular, when $\mathcal{L}=\mathcal{T}_{X}$ over a non-singular $a$-space (smooth manifold, complex manifold, smooth algebraic variety or smooth scheme over the field $\mathbb{C})$ $X$, we get isomorphism of graded vector spaces
$\mathbb{H}H^{\bullet}(\mathscr{J}(\mathcal{O}_{X},\mathcal{T}_{X}))\cong\mathbb{H}^{\bullet}(X,\Omega^{\bullet}_{X})\cong H^{\bullet}(X,\mathbb{K})$ or $H^{\bullet}(X^{an},\mathbb{K})$
by applying de Rham theorems in different settings ((((smooth, analytic, algebraic)))) [9, 36], where $\mathbb{K=R}$ or $\mathbb{C}$ accordingly and $X^{an}$ is the analytification of the algebraic variety (scheme) $X$.
Corollary 5.15.
Applying the Theorem 3.6 for $\mathcal{L}=\mathcal{T}_{X}$ over some non-singular $a$-space $X$ and using the Corollary 5.14 we get the isomorphism
$\mathbb{H}H^{\bullet}(\mathscr{J}(\mathcal{O}_{X},\mathcal{T}_{X}))\cong Ext^{\bullet}_{\mathcal{D}_{X}}(\mathcal{O}_{X},\mathcal{O}_{X}),$
where $\mathcal{D}_{X}:=\mathcal{D}iff(\mathcal{O}_{X})\cong\mathscr{U}(\mathcal{O}_{X},\mathcal{T}_{X})$ is the sheaf of differential operators of $X$.
Thus, in this case we get a canonical isomorphism
$Cotor^{\bullet}_{\mathcal{J}_{X}}(\mathcal{O}_{X},\mathcal{O}_{X})\cong Ext^{\bullet}_{\mathcal{D}_{X}}(\mathcal{O}_{X},\mathcal{O}_{X}),$
where $\mathcal{J}_{X}:=\mathscr{J}(\mathcal{O}_{X},\mathcal{T}_{X})$ is the usual sheaf of jets on $X$ [5, 32].
5.4. Remarks on Lie algebroids over Schemes.
All the previous discussions are applicable to Lie algebroids over noetherian separated schemes (or schemes of finite types) over a field of characteristic zero.